### v1

# Symplectic singularities arising from algebras of symmetric tensors

Abstract.
The algebra of symmetric tensors $S(X)\coloneqq H^{0}(X,{\sf S}^{\bullet}T_{X})$ of a projective manifold $X$ leads to a natural dominant affinization morphism
$\varphi_{X}\colon T^{*}X\longrightarrow\mathcal{Z}_{X}\coloneqq\operatorname{Spec}S(X).$
It is shown that $\varphi_{X}$ is birational if and only if $T_{X}$ is big. We prove that if $\varphi_{X}$ is birational, then $\mathcal{Z}_{X}$ is a symplectic variety endowed with the Schouten–Nijenhuis bracket if and only if $\mathbb{P}T_{X}$ is of Fano type, which is the case for smooth projective toric varieties, smooth horospherical varieties with small boundary and the quintic del Pezzo threefold. These give examples of a distinguished class of conical symplectic varieties, which we call symplectic orbifold cones.
Key words and phrases:
symplectic singularity, cotangent bundle, hypertoric variety, horospherical variety, Fano variety, moment map
2020 Mathematics Subject Classification:
14B05, 14J42, 14J45, 14M17, 14M25
1. Introduction
1.1. The algebra of symmetric tensors
Let $X$ be a projective manifold and let $T^{*}X$ be the total space of the cotangent bundle of $X$. Denote by $\eta$ the Liouville $1$-form on $T^{*}X$. Then $T^{*}X$ carries a natural holomorphic symplectic form $\omega=d\eta$, rendering it a non-compact holomorphic symplectic manifold. Although the symplectic geometry of these manifolds has been studied extensively in the literature for some (very limited) special examples (see Example 1.1 and Example 1.2 below), the general case still remains intriguing [Hwa15].
Recall that the algebra of symmetric tensors $S(X)$ is the graded $\mathbb{C}$-algebra defined as follows:
$S(X)\coloneqq\bigoplus_{p\geq 0}H^{0}(X,{\sf S}^{p}T_{X})=\bigoplus_{p\geq 0}H^{0}(\mathbb{P}T_{X},\scr{O}_{\mathbb{P}T_{X}}(p))=H^{0}(T^{*}X,\scr{O}_{T^{*}X}).$
Then $S(X)$ is an integrally closed domain and the symplectic structure over $T^{*}X$ defines a Poisson structure over $S(X)$, called the Schouten–Nijenhuis bracket. Denote by $\mathcal{Z}_{X}$ the normal integral affine scheme $\operatorname{Spec}S(X)$. Then we have the natural dominant affinization morphism
$\varphi_{X}\colon T^{*}X\longrightarrow\mathcal{Z}_{X}.$
Despite its simple definition, the affine scheme $\mathcal{Z}_{X}$ together with the morphism $\varphi_{X}$ has been little studied (see [BL24]). It turns out that $S(X)=\mathbb{C}$ in many cases and hence $\varphi_{X}$ is a constant morphism ([BL24, § 3]). However, one should bear in mind that in general $\varphi_{X}$ is a dominant but non-surjective morphism.
In the following two examples, we collect some samples of interesting interactions between the symplectic geometry of cotangent bundles and the map $\varphi_{X}$.
Example 1.1 (Hitchin map).
Let $C$ be a smooth projective curve with genus $g\geq 2$. We fix co-prime integers $r$, $d\geq 1$ and a line bundle $L$ of degree $d$ on $C$. Consider the moduli space $\mathcal{M}_{C}(r,L)$ of stable vector bundles $E$ on $C$ with rank $r$ and $\det(E)\cong L$. Then $\mathcal{M}_{C}(r,L)$ is a smooth projective variety. By deformation theory, the cotangent bundle of $\mathcal{M}_{C}(r,L)$ at a point $[E]$ is isomorphic to $\operatorname{Hom}_{0}(E,E\otimes K_{C})$, i.e., the traceless $K_{C}$-valued endomorphisms of $E$. Then there exists a surjective morphism, called the Hitchin map, defined as follows:
${\sf h}_{C}\colon T^{*}\mathcal{M}_{C}(r,L)$ $\longrightarrow\bigoplus_{i=2}^{r}H^{0}(C,K_{C}^{i})\eqqcolon\mathcal{A}_{C}$
$(E,u)$ $\longmapsto(\textup{Tr}\wedge^{2}u,\dots,\textup{Tr}\wedge^{r}u).$
It was proved in [Hit87] that ${\sf h}_{C}$ is an algebraically completely integrable system; that is,
  1. (1.1.1)
    the map ${\sf h}_{C}$ is surjective and Lagrangian, and
  2. (1.1.2)
    the general fiber ${\sf h}_{C}^{-1}(a)$ is of the form $A\setminus Z$, where $A$ is an Abelian variety and $Z$ is a closed subset with co-dimension $\geq 2$.
Consequently, the Hitchin map ${\sf h}_{C}$ coincides with $\varphi_{\mathcal{M}_{C}(r,L)}$ and $\mathcal{Z}_{\mathcal{M}_{C}(r,L)}$ is isomorphic to the $(r^{2}-1)(g-1)$-dimensional affine space $\mathcal{A}_{C}$ ([BL24, § 2.7]). A similar phenomenon also happens for smooth complete intersections of two quadratic hypersurfaces in $\mathbb{P}^{n+2}$ for $n\geq 2$ [BEH+23].
Example 1.2 (Springer map).
Let $G$ be a semi-simple linear algebraic group and let $P\subset G$ be a parabolic subgroup. We denote by $\mathfrak{g}$ and $\mathfrak{p}$ their Lie algebras. Then the cotangent bundle $T^{*}(G/P)$ of the rational homogeneous space $G/P$ can be naturally identified to the homogeneous bundle $G\times_{P}\mathfrak{p}^{\perp}$. Associating to a pair $(g,v)\in G\times\mathfrak{p}^{\perp}$ the element $\textup{Ad}(g)\cdot v$ of $\mathfrak{g}^{*}$ defines a generically finite proper map $\Phi\colon T^{*}(G/P)\rightarrow\mathfrak{g}^{*}$, called the Springer map, whose image $\widebar{\mathcal{O}}_{P}$ is the closure of a nilpotent orbit [Ric74]. Let
$T^{*}(G/P)\stackrel{{\scriptstyle\widetilde{\Phi}}}{{\longrightarrow}}\widetilde{\mathcal{O}}_{P}\longrightarrow\widebar{\mathcal{O}}_{P}$
be the Stein factorization of $\Phi$. Then $\widetilde{\Phi}$ is a birational projective morphism, and hence $\widetilde{\mathcal{O}}_{P}=\mathcal{Z}_{G/P}$ and $\varphi_{G/P}=\widetilde{\Phi}$.
1.2. Symplectic varieties
Recall that a normal variety $Z$ is symplectic in the sense of Beauville [Bea00] if there exists a symplectic form $\omega$ on the smooth locus $Z_{\operatorname{reg}}$ of $Z$ such that for any resolution $W\rightarrow Z$, the $2$-form $\omega$ extends to a holomorphic $2$-form on $W$. Moreover, a resolution $W\rightarrow Z$ of a symplectic variety $Z$ is called a symplectic resolution if $\omega$ extends to a symplectic form on $W$. In Example 1.2 above, the natural symplectic form on $T^{*}(G/P)$ induces a symplectic form on $\widetilde{\mathcal{O}}_{P,\operatorname{reg}}$ such that the affine normal variety $\widetilde{\mathcal{O}}_{P}$ is symplectic and $\widetilde{\Phi}$ is a symplectic resolution. We refer the reader to [Fu06] for more details on symplectic singularities.
The main goal of this paper is to pursue further within the spirit of Example 1.2, studying the symplectic geometry of cotangent bundles of projective manifolds from the viewpoint of symplectic singularities. To this end, we introduce the notion of symplectic orbifold cones and we refer the reader to Definition 2.8 and Remark 3.6 for the justification of this terminology.
Definition 1.3.
Let $X$ be a projective manifold. We call $\mathcal{Z}_{X}$ a symplectic orbifold cone if $\varphi_{X}$ is birational and $\mathcal{Z}_{X}$ is a symplectic variety endowed with the Schouten–Nijenhuis bracket.
We are interested in the following questions inspired by Example 1.2.
Question 1.4.
Let $X$ be a projective manifold and $\varphi_{X}\colon T^{*}X\rightarrow\mathcal{Z}_{X}$ the associated morphism.
  1. (1.4.1)
    When is $\varphi_{X}$ birational? When is $\mathcal{Z}_{X}$ a symplectic variety endowed with the Schouten–Nijenhuis bracket in the case that $\varphi_{X}$ is birational?
  2. (1.4.2)
    Characterize symplectic orbifold cones among symplectic varieties, determine their possible symplectic resolutions and study its relation with the geometry of $X$ itself.
Question 1.4 is also closely related to the following questions, which are of independent interests.
Question 1.5.
Let $X$ be a projective manifold.
  1. (1.5.1)
    (Beauville) Is the $\mathbb{C}$-algebra $S(X)$ finitely generated?
  2. (1.5.2)
    (Campana–Peternell) If $\varphi_{X}\colon T^{*}X\rightarrow\mathcal{Z}_{X}$ is a symplectic resolution, is $X$ a rational homogeneous space?
It turns out that Question (1.5.2) is equivalent to the Campana–Peternell conjecture, which asserts that a simply connected projective manifold has nef tangent bundle if and only if it is a rational homogeneous space (see Lemma 2.7 and [Wan24])
1.3. Main results
Recall that a normal projective variety $X$ is said to be of Fano type if there exists an effective $\mathbb{Q}$-divisor $\Delta$ such that $-(K_{X}+\Delta)$ is ample and the pair $(X,\Delta)$ is klt. A vector bundle $E$ over a projective manifold is called big if $\scr{O}_{\mathbb{P}E}(1)$ is so, where $\mathbb{P}E$ is the projectivization of $E$ in the sense of Grothendieck.
Theorem 1.6.
Let $X$ be a projective manifold.
  1. (1.6.1)
    The map $\varphi_{X}$ is birational if and only if $T_{X}$ is big.
  2. (1.6.2)
    The affine scheme $\mathcal{Z}_{X}$ is a symplectic orbifold cone if and only if $T_{X}$ is big and $\mathbb{P}T_{X}$ is of Fano type. In particular, if $\mathcal{Z}_{X}$ is a symplectic orbifold cone, then $X$ itself is of Fano type.
The bigness of $T_{X}$ is a rather restrictive property for projective manifolds (see for instance [HLS22, KKL22, HL23, SZ23] and the references therein). The known interesting examples of projective manifolds with big tangent bundles include the quintic del Pezzo threefold [HLS22] and smooth projective horospherical varieties ([Liu23, § 3B2]), which contain toric varieties [Hsi15] and rational homogeneous spaces [Ric74]. Applying Theorem 1.6 to these examples yields the following list of examples of symplectic orbifold cones.
Theorem 1.7.
The affine scheme $\mathcal{Z}_{X}$ is a symplectic orbifold cone if $X$ is isomorphic to one of the following:
  1. (1.7.1)
    a smooth projective toric variety;
  2. (1.7.2)
    a smooth projective horospherical variety with small boundary;
  3. (1.7.3)
    the quintic del Pezzo threefold.
It is unclear for us whether the statement still holds for all smooth projective horospherical varieties. Each of the three cases requires a different proof, which will be discussed case-by-case in the following.
1.3.1. Toric case
For a smooth projective toric variety $X$ defined by a fan $\Sigma$, we have the natural exact sequence
$0\longrightarrow\mathbf{M}\stackrel{{\scriptstyle B}}{{\longrightarrow}}\mathbb{Z}^{\Sigma(1)}=\mathbb{Z}^{N}\stackrel{{\scriptstyle A}}{{\longrightarrow}}\operatorname{Pic}(X)\longrightarrow 0,$
where $\Sigma(1)$ is the set of one-dimensional rays in $\Sigma$ and $\mathbf{M}$ is the lattice of characters of the torus. By using the integer matrix $A$, we can associate $X$ with a toric hyperKähler variety, denoted by $\mathcal{Y}(A,0,0)$ (§ 4.2), which is known to be a symplectic variety by [Bel23]. Based on an explicit computation of $S(X)$, it turns out that $\mathcal{Z}_{X}$ coincides with $\mathcal{Y}(A,0,0)$, namely we have
Theorem 1.8.
Let $X$ be a smooth projective toric variety. Then $\mathcal{Z}_{X}$ is isomorphic to the toric hyperKähler variety $\mathcal{Y}(A,0,0)$.
Note that the existence problem of symplectic resolutions for toric hyperKähler varieties is completely solved by Y. Namikawa in [Nam23, Corollary 13] and thus we get a satisfactory answer to Question 1.4 for toric varieties. Moreover, as an immediate application of Theorem 1.6 and Theorem 1.8, we obtain the following seemly-new result, which generalizes [HS10, Theorem 5.9].
Corollary 1.9.
The projectivized tangent bundle $\mathbb{P}T_{X}$ of a smooth projective toric variety $X$ is of Fano type.
1.3.2. Horospherical case
Recall that a $G$-variety is said to be horospherical if it contains an open $G$-orbit which is a torus bundle over a projective rational homogeneous space (§ 5.1). Toric varieties are examples of horospherical varieties. For a horospherical variety $X$, we will relate the Cox ring of $\mathbb{P}T_{X}$ to that of a weak Fano manifold and then a characterization of varieties of Fano type via Cox rings obtained in [GOST15] allows us to obtain the following result.
Theorem 1.10.
Let $X$ be a smooth projective $G$-horospherical variety. Then $X$ is of Fano type and $\mathbb{P}T_{X}$ is a Mori dream space. If additionally $X$ does not contain any $G$-stable divisor, then $\mathbb{P}T_{X}$ is also of Fano type.
According to the classification result obtained by B. Pasquier in [Pas09], any smooth projective horopsherical $G$-variety $X$ with Picard number one contains no $G$-stable divisors and thus $\mathcal{Z}_{X}$ is a symplectic variety. For such $X$, the identity component of its automorphism group $\widetilde{G}=\operatorname{Aut}^{0}(X)$ is worked out in loc. cit.. In particular, the variety $X$ carries a natural action by a torus $\mathbb{T}\cong\mathbb{C}^{\times}$, which is transitive and free along the torus fibers of the horospherical open orbit, such that $\widebar{G}\coloneqq G\times\mathbb{T}$ is a subgroup of $\widetilde{G}$. We consider the associated moment maps (§ 2.4) for the actions of $\widebar{G}$ and $\widetilde{G}$, respectively, on $T^{*}X$:
[…]
Then we can relate $\mathcal{Z}_{X}$ to the moment map images $\mathcal{M}_{X}^{\widetilde{G}}$ and $\mathcal{M}_{X}^{\widebar{G}}$. Note that the maps $\Phi_{X}^{\widetilde{G}}$ and $\Phi_{X}^{\widebar{G}}$ factor through $\varphi_{X}\colon T^{*}X\rightarrow\mathcal{Z}_{X}$ via the natural maps ${\sf S}^{\bullet}\widetilde{\mathfrak{g}}\rightarrow S(X)$ and ${\sf S}^{\bullet}\widebar{\mathfrak{g}}\rightarrow S(X)$.
Proposition 1.11.
Let $X$ be a smooth projective horospherical $G$-variety with Picard number one, and let $\widetilde{G}$ be the identify component of its automorphism group. Then the induced map $\mathcal{Z}_{X}\rightarrow\mathcal{M}_{X}^{\widetilde{G}}$ is a generically finite morphism satisfying the following commutative diagram
[…]
where ~$\widetilde{\mathcal{M}}_{X}\cong\mathcal{Z}_{X}/\!\!/\mathbb{T}$, $\mathcal{M}_{X}^{\widebar{G}}\cong\mathcal{M}_{X}^{\widetilde{G}}/\!\!/\mathbb{T}$ and ~$\widetilde{\mathcal{M}}_{X}\rightarrow\mathcal{M}_{X}^{\widebar{G}}$ is a finite morphism.
1.3.3. The quintic del Pezzo 3-fold
The quintic del Pezzo $3$-fold (denoted by $V_{5}$) is a smooth codimension $3$ linear section of ${\rm Gr}(2,5)$ in the Plücker embedding. Recently A. Höring and T. Peternell have constructed explicitly a weak Fano model of $\mathbb{P}T_{V_{5}}$ in [HP24], which in particular implies that $\mathbb{P}T_{V_{5}}$ is of Fano type, and hence $\mathcal{Z}_{V_{5}}$ is actually a $6$-dimensional symplectic variety by Theorem 1.6. Following the explicit construction in [HP24], we prove that the singular locus $Z$ of $\mathcal{Z}_{V_{5}}$ is a codimension two irreducible closed subset, the transverse singularity of $\mathcal{Z}_{V_{5}}$ along a general point of $Z$ is an $A_{2}$-singularity, and $\mathcal{Z}_{V_{5}}$ admits a unique $\mathbb{Q}$-factorial terminalization, which can be constructed explicitly and gives a symplectic resolution of $\mathcal{Z}_{V_{5}}$ outside the vertex (Proposition 6.10). In particular, this allows us to derive that $\mathcal{Z}_{V_{5}}$ does not admit any symplectic resolution (Proposition 6.11).
2. Preliminaries
Throughout this paper, we work over the field $\mathbb{C}$ of complex numbers. Let $\mathbb{T}$ denote the $1$-dimensional torus $\mathbb{C}^{\times}$. All varieties and manifolds are always assumed to be irreducible. We will always denote by $\pi\colon T^{*}X\rightarrow X$ and $\widebar{\pi}\colon\mathbb{P}T_{X}\rightarrow X$ the natural projections, respectively.
2.1. Contact varieties
Recall that a contact manifold $(X,L)$ is a pair consisting of a nonsingular variety $X$, a Weil divisor $L$ on $X$ and an exact sequence of vector bundles
[…]
such that the induced map by Lie bracket
[…]
is non-degenerate at every point of $X$; or equivalently the twisted form
$\theta\wedge(d\theta)^{n-1}\in H^{0}(X,K_{X}+nL)$ (2.0.1)
is nonwhere vanishing, where $\dim(X)=2n-1$. In particular, we have $K_{X}+nL\sim 0$. We call $\scr{O}_{X}(L)$ the contact line bundle and the $\scr{O}_{X}(L)$-valued $1$-form $\theta$, viewed as a global section of $\scr{O}_{X}(L)\otimes\Omega_{X}^{1}$, is called the contact form. Finally we remark that the degeneracy locus of $\theta$ has pure codimension one and hence the non-degeneracy of $\theta$ can be verified over any open subset $U\subset X$ with $\operatorname{codim}(X\setminus U)\geq 2$.
Example 2.1 ([KPSW00, § 2.6]).
Let $X$ be a smooth variety. Then the projectivized tangent bundle $(\mathbb{P}T_{X},\scr{O}_{\mathbb{P}T_{X}}(1))$ is a contact manifold. The contact form $\theta$ is given by the following composition:
$T_{\mathbb{P}T_{X}}\longrightarrow\widebar{\pi}^{*}T_{X}\rightarrow\scr{O}_{\mathbb{P}T_{X}}(1).$
More generally, a quasi-contact variety $(X,L)$ is a pair consisting of a normal variety $X$ and a Weil divisor $L$ on $X$ such that $(X_{\operatorname{reg}},L|_{X_{\operatorname{reg}}})$ is a contact manifold and a quasi-contact variety $(X,L)$ is called contact if $L$ is Cartier.
2.2. Orbifold cones
An affine variety $X=\operatorname{Spec}R$ is called a cone if there exists a positive grading $R=\oplus_{i\geq 0}R_{i}$ such that $R_{0}=\mathbb{C}$ and $R$ is a finitely generated domain over $R_{0}$. The grading determines a natural $\mathbb{T}$-action on $X$ and the origin $0\in X$ defined by the maximal ideal $\mathfrak{m}\coloneqq\oplus_{i\geq 1}R_{i}$ is the unique fixed point of this $\mathbb{C}^{\times}$-action.
Example 2.2.
Let $X$ be a normal variety and let $L$ be a $\mathbb{Q}$-Cartier Weil divisor on $X$. The relative spectrum $\mathcal{L}$ and the section ring $R(L)$ are defined as follows:
$\mathcal{L}\coloneqq\operatorname{Spec}_{X}\left(\bigoplus_{p\geq 0}\scr{O}_{X}(pL)\right)\quad\textup{and}\quad R(L)\coloneqq\bigoplus_{p\geq 0}H^{0}(X,\scr{O}_{X}(pL)).$
Then the natural projection $p\colon\mathcal{L}\rightarrow X$ is an affine morphism [ADHL15, Proposition 1.3.2.3]. We also denote by $\mathcal{L}^{\times}$ the associated Seifert $\mathbb{C}^{\times}$-bundle [Kol04]:
×$\mathcal{L}^{\times}\coloneqq\operatorname{Spec}_{X}\left(\bigoplus_{p\in\mathbb{Z}}\scr{O}_{X}(pL)\right).$
Then $\mathcal{L}^{\times}$ is the open subset of $\mathcal{L}$ obtained by removing the zero section $\textbf{0}_{\mathcal{L}}$ of $\mathcal{L}$ [Kol04, 14]. Let $\mathcal{L}_{\operatorname{reg}}$ be the preimage $p^{-1}(X_{\operatorname{reg}})$. Then $\mathcal{L}_{\operatorname{reg}}$ is actually the total space of the line bundle $\scr{O}_{X_{\operatorname{reg}}}(-L)$ over $X_{\operatorname{reg}}$ and we observe that the natural isomorphism $R(L)\rightarrow R(L|_{X_{\operatorname{reg}}})$ yields an identification of $R(L)$ to the ring of regular functions over $\mathcal{L}_{\operatorname{reg}}$. So $R(L)$ is normal. We will denote by $\mathcal{C}_{L}$ the normal affine scheme $\operatorname{Spec}R(L)$ and by $\mathbb{P}(\mathcal{C}_{L})$ the projective scheme $\operatorname{Proj}R(L)$. Then there exists a natural dominant affinization morphism
$\varphi_{L}\colon\mathcal{L}\longrightarrow\mathcal{C}_{L}$
and $\mathcal{C}_{L}$ is a cone if $H^{0}(X,\scr{O}_{X})=\mathbb{C}$ and $R(L)$ is a finitely generated $\mathbb{C}$-algebra.
Lemma 2.3.
Let $X$ be a normal projective variety and let $L$ be a $\mathbb{Q}$-Cartier Weil divisor on $X$ such that $R(L)$ is finitely generated. Then $\varphi_{L}$ is proper if and only if $L$ is semi-ample.
Proof.
First we assume that $L$ is Cartier. Then $\mathcal{L}\rightarrow X$ is a $\mathbb{C}$-bundle. If $\varphi_{L}$ is proper, it follows that $\varphi_{L}^{-1}(0)$ is complete and thus $\varphi_{L}^{-1}(0)=\textbf{0}_{\mathcal{L}}$ since $\varphi_{L}$ is $\mathbb{C}^{\times}$-equivariant. This means exactly that $L$ is semi-ample. On the other hand, suppose that $L$ is semi-ample and consider the contraction $\widebar{\varphi}_{L}\colon X\rightarrow\mathbb{P}(\mathcal{C}_{L})$. Then there exists an ample Cartier divisor $\widebar{L}$ on $\mathbb{P}(\mathcal{C}_{L})$ such that $\widebar{\varphi}_{L}^{*}\widebar{L}=L$ and $\widebar{\varphi}_{L}^{*}\colon R(\widebar{L})\rightarrow R(L)$ is an isomorphism of graded $\mathbb{C}$-algebras. Moreover, as $\mathcal{L}=X\times_{\mathbb{P}(\mathcal{C}_{L})}\widebar{\mathcal{L}}$, the map $\varphi_{L}$ is proper if and only if $\varphi_{\widebar{L}}$ is proper. The latter is true since $\varphi_{\widebar{L}}\colon\widebar{\mathcal{L}}\rightarrow\mathcal{C}_{\widebar{L}}=\mathcal{C}_{L}$ is actually a weighted blow-up of $0\in\mathcal{C}_{L}$ (cf. [Kol13, § 3.1])
Next we choose a positive integer $m$ such that $L^{\prime}\coloneqq mL$ is Cartier and consider the following commutative diagram
${\mathcal{L}}$${\mathcal{C}_{L}}$${\mathcal{L}^{\prime}}$${\mathcal{C}_{L^{\prime}}}$$\scriptstyle{\varphi_{L}}$$\scriptstyle{q^{\prime}}$$\scriptstyle{q}$$\scriptstyle{\varphi_{L^{\prime}}}$
Note that $q^{\prime}$ and $q$ are finite morphisms (see [ADHL15, Proposition 1.1.2.5] and its proof), so $\varphi_{L}$ is proper if and only if $\varphi_{L^{\prime}}$ is proper. Then the result follows because $L$ is semi-ample if and only if $L^{\prime}$ is semi-ample. ∎
Lemma 2.4.
Let $X$ be a normal projective variety and let $L$ be a $\mathbb{Q}$-Cartier Weil divisor on $X$. Then the affinization morphism $\varphi_{L}\colon\mathcal{L}\rightarrow\mathcal{C}_{L}$ is birational if and only if $L$ is big.
Proof.
By [Cut14, Lemma 7.2], $L$ is big if and only if $\dim R(X)=\dim(X)+1$. In particular, as $\dim R(L)=\dim\mathcal{C}_{L}$, it follows that $L$ is big if $\varphi_{L}$ is birational. Now we suppose that $L$ is big and consider the following commutative diagram
[…]
Since $\widebar{\varphi}_{L}\colon X\dashrightarrow\mathbb{P}(\mathcal{C}_{L})$ is birational and $\varphi_{L}$ is $\mathbb{C}^{\times}$-equivariant, it remains to show that for a general point $x\in X$, the restriction of $\varphi_{L}$ to the fiber $\mathbb{C}\cong\mathcal{L}_{x}\coloneqq p^{-1}(x)$ is generically injective. Recall that there exists some $m\gg 1$ such that $H^{0}(X,\scr{O}_{X}(mL))\not=0$ and $H^{0}(X,\scr{O}_{X}((m+1)L))\not=0$ ([Laz04, Corollary 2.2.10]). Let $s_{i}\in H^{0}(X,\scr{O}_{X}(iL))$ be a non-zero element for $i=m$ and $m+1$. As $x$ is general, we may also assume that $s_{i}|_{\mathcal{L}_{x}}\not\equiv 0$. For two non-zero points $v$ and $v^{\prime}\in\mathcal{L}_{x}$, there exists a constant $\lambda\in\mathbb{C}^{\times}$ such that $v^{\prime}=\lambda v$. In particular, we have $s_{i}(v^{\prime})=\lambda^{i}s_{i}(v)$. As $(m,m+1)=1$, it follows that $s_{i}(v)=s_{i}(v^{\prime})$, $i\in\{m,m+1\}$, if and only if $\lambda=1$, i.e., $v=v^{\prime}$. ∎
Corollary 2.5.
Let $X$ be a projective manifold.
  1. (2.5.1)
    The map $\varphi_{X}:T^{*}X\to\mathcal{Z}_{X}$ is proper if and only if $T_{X}$ is semi-ample.
  2. (2.5.2)
    The map $\varphi_{X}$ is birational if and only if $T_{X}$ is big.
Proof.
As in Example 2.1, we consider the contact manifold $(\mathbb{P}T_{X},L)$, where $L$ is a divisor corresponding to $\scr{O}_{\mathbb{P}T_{X}}(1)$. Then we have the following commutative diagram:
[…]
where $\mu$ is the blow-up along the zero section $\textbf{0}_{X}$ of $T^{*}X\rightarrow X$. The result follows since $T_{X}$ is semi-ample (resp. big) if and only if $L$ is semi-ample (resp. big). ∎
As an application, one can relate the Campana–Peternell’s question (1.5.2) to the following weak version of Campana–Peternell’s conjecture, which is equivalent to the Campana–Peternell conjecture by a very recent result of J. Wang [Wan24].
Conjecture 2.6 ([CP91, 11.2]).
Let $X$ be a projective manifold such that $T_{X}$ is big and nef. Then $X$ is a rational homogeneous space.
Lemma 2.7.
Conjecture 2.6 holds if and only if the answer to Question (1.5.2) is positive.
Proof.
By Corollary 2.5, the map $\varphi_{X}\colon T^{*}X\rightarrow\mathcal{Z}_{X}$ is a symplectic resolution if and only if $T_{X}$ is big and semi-ample. On the other hand, by the Basepoint Free Theorem, the latter is equivalent to that $T_{X}$ is big and nef. ∎
One of the most important examples of cones is the orbifold cones, which is a generalization of cones over Fano manifolds — see [Zhu23, § 3.1] and [Kol13, § 3.1] for more details.
Definition 2.8.
Let $X$ be a normal projective variety and let $L$ be an ample $\mathbb{Q}$-Cartier Weil divisor on $X$. Then the affine cone $\mathcal{C}_{L}$ is called an orbifold cone if $-K_{X}\sim_{\mathbb{Q}}rL$ for some $r\in\mathbb{Q}_{>0}$.
Lemma 2.9 ([Kol04, 42] and [Zhu23, Lemma 3.3]).
An orbifold cone $\mathcal{C}_{L}$ is klt if and only if $X$ itself is klt.
2.3. Conical symplectic varieties
A symplectic variety $(Z,\omega)$ is called conical if $Z=\operatorname{Spec}(R)$ is an affine cone and $\omega$ is homogeneous with respect to the natural $\mathbb{T}$-action; that is, there exists a non-zero integer $l\in\mathbb{Z}$ such that $\lambda^{*}\omega=\lambda^{l}\omega$ for any $\lambda\in\mathbb{T}$. In this case we call $l$ the weight of $(Z,\omega)$. Symplectic orbifold cones form a particularly interesting class of conical symplectic varieties with weight $1$. An interesting question is to characterize symplectic orbifold cones among all conical symplectic varieties. Now we study symplectic resolutions of conical symplectic varieties. The following result is from [Fu03, Lemmas 3.5-3.7]:
Proposition 2.10.
Assume that $(Z,\omega)$ is a conical symplectic variety with weight $1$ and $\mu\colon Y\to Z$ is a symplectic resolution with the lifted symplectic structure $\widetilde{\omega}$. Then there exists an irreducible component $F$ of $\mu^{-1}(0)$, and an open subset $U\subset Y$ such that $(U,\widetilde{\omega})\simeq(T^{*}F,\omega_{s})$ as conical symplectic varieties, where $\omega_{s}$ is the standard symplectic structure on the cotangent bundle.
Corollary 2.11.
Under the same assumption as in Proposition 2.10, if moreover $\mu^{-1}(0)$ is irreducible, then $Y=T^{*}F$. If additionally $\dim Z\leq 10$, then $F$ is a rational homogeneous space and $\mu$ is a Springer map.
Proof.
As $\mu^{-1}(0)$ is irreducible, we have $\mu^{-1}(0)=F$, hence $Y^{\mathbb{T}}=F$. This implies that $Y\simeq T^{*}F$. It then follows that $Z\simeq\mathcal{Z}_{F}$ and $T_{F}$ is nef and big. By [Kan17], the variety $F$ is a rational homogeneous space if $\dim F\leq 5$, which concludes the proof. ∎
2.4. Moment maps
Let $G$ be a linear algebraic group and denote by $\mathfrak{g}$ its Lie algebra. Let $X$ be a smooth variety equipped with a $G$-action. Given an arbitrary point $x\in X$, the tangent map of the orbit map $\mu_{x}\colon G\rightarrow Gx$, $g\mapsto gx$, defines a linear map
$d\mu_{x}|_{e}\colon\mathfrak{g}\longrightarrow T_{x}(Gx)\subset T_{x}X.$
Taking dual yields a linear map $T^{*}_{x}X\rightarrow\mathfrak{g}^{*}$ and thus a morphism $\Phi_{X}^{G}\colon T^{*}X\rightarrow\mathfrak{g}^{*}$, which is called the moment map (associated to the $G$-action). By $\mathcal{M}_{X}^{G}\subset\mathfrak{g}^{*}$ we denote the closure of its image. Denote the image of the natural map ${\sf S}^{\bullet}\mathfrak{g}\rightarrow S(X)$ by $R$. Then $\mathcal{M}_{X}^{G}\cong\operatorname{Spec}R$ and we have a natural commutative diagram
[…]
The projectivization of $\Phi_{X}^{G}$ by the natural $\mathbb{T}$-action induces a rational map $\mathbb{P}T_{X}\dashrightarrow\mathbb{P}(\mathcal{M}_{X}^{G})$, which is exactly the map defined by the linear system induced by the linear subspace $\mathfrak{g}\subset H^{0}(X,T_{X})\cong H^{0}(\mathbb{P}T_{X},\scr{O}_{\mathbb{P}T_{X}}(1))$ ([Liu23, § 2C]).
Example 2.12.
Let $X=\mathbb{C}^{n}$ be the $n$-dimensional complex affine space. Then we can identify $T^{*}X$ to $\mathbb{C}^{n}\times(\mathbb{C}^{n})^{*}$ with coordinates $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$.
  1. (2.12.1)
    Assume that the torus $\mathbb{T}$ acts on $\mathbb{C}^{n}$ with weight $(a_{1},\dots,a_{n})$; that is,
    $t\cdot(x_{1},\dots,x_{n})\coloneqq(t^{a_{1}}x_{1},\dots,t^{a_{n}}x_{n}).$
    Then the moment map $\Phi_{X}^{\mathbb{T}}$ can be written in coordinates as follows:
    $\mathbb{C}^{n}\times(\mathbb{C}^{n})^{*}\rightarrow\mathfrak{t}^{*}\cong\mathbb{C},\quad(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\mapsto\sum_{i=1}^{n}a_{i}x_{i}y_{i}.$
  2. (2.12.2)
    Assume that the vector group $\mathbb{V}\coloneqq\mathbb{C}^{n}$ acts on $\mathbb{C}^{n}$ by translation; that is,
    $(v_{1},\dots,v_{n})\cdot(x_{1},\dots,x_{n})\coloneqq(v_{1}+x_{1},\dots,v_{n}+x_{n}).$
    Then the moment map $\Phi_{X}^{\mathbb{V}}$ is nothing but the second projection, i.e.,
    $\mathbb{C}^{n}\times(\mathbb{C}^{n})^{*}\rightarrow\mathfrak{v}^{*}\cong\mathbb{C}^{n},\quad(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\mapsto(y_{1},\dots,y_{n}).$
Here are some properties of the moment map:
Proposition 2.13.
Let $G$ be a linear algebraic group and let $X$ be a smooth projective $G$-variety.
  1. (2.13.1)
    The moment map $\Phi_{X}^{G}$ is generically finite if and only if $G$ acts on $T^{*}X$ with an open orbit. In this case, the image $\mathcal{M}_{X}^{G}$ is the closure of a co-adjoint orbit in $\mathfrak{g}^{*}$.
  2. (2.13.2)
    The moment map $\Phi_{X}^{G}$ is proper if and only if $X$ is a projective $G$-homogeneous variety.
  3. (2.13.3)
    Assume that $G$ is reductive. Then $\Phi_{X}^{G}$ is generically finite if and only if $X$ is isomorphic to $G/P$ and the map $\Phi_{X}^{G}$ is the Springer map.
Proof.
It is well-known (cf. [Tim11, § 8.2]) that for a point $\alpha\in T^{*}X$, the kernel ${\rm Ker}d_{\alpha}\Phi_{X}^{G}$ is the skew-orthocomplement of the tangent space of $G$-orbit $G\cdot\alpha$ and the image $d_{\alpha}\Phi_{X}^{G}$ is the annihilator of the stabilizer $\mathfrak{g}_{\alpha}$. It follows that $\Phi_{X}^{G}$ is generically finite if and only if $G$ acts on $T^{*}X$ with an open orbit. In this case, the image $\mathcal{M}_{X}^{G}$ contains an open $G$-orbit, hence it is the closure of the co-adjoint orbit. This proves (2.13.1).
For (2.13.2), assume that $\Phi_{X}^{G}$ is proper, then $(\Phi_{X}^{G})^{-1}(0)$ is complete. By the $\mathbb{C}^{\times}$-equivariance, we have $X\subseteq(\Phi_{X}^{G})^{-1}(0)$, which we claim to be an equality. Assume $(x,\xi)\in T^{*}X$ is contained in $(\Phi_{X}^{G})^{-1}(0)$ for some $0\neq\xi\in T_{x}^{*}X$, then $(x,\lambda\xi)$ is also contained in $(\Phi_{X}^{G})^{-1}(0)$ for all $\lambda\in\mathbb{C}^{\times}$. As $(\Phi_{X}^{G})^{-1}(0)$ is complete, this implies that $\lim_{\lambda\to\infty}(x,\lambda\xi)$ lies in $T^{*}X$, which is absurd. This proves that $X=(\Phi_{X}^{G})^{-1}(0)$, hence we have a morphism $\widebar{\Phi}\colon\mathbb{P}T_{X}\to\mathbb{P}\mathfrak{g}^{*}$ such that $\scr{O}_{\mathbb{P}T_{X}}(1)=\widebar{\Phi}^{*}(\scr{O}_{\mathbb{P}\mathfrak{g}^{*}}(1))$ is globally generated, hence $T_{X}$ is globally generated by $\mathfrak{g}\subset H^{0}(X,T_{X})$, which implies that $X$ is $G$-homogeneous.
For (2.13.3), by [Tim11, Theorem 8.17], the map $\Phi_{X}^{G}$ is generically finite if and only if the $G$-action on $X$ has complexity 00 and rank 00, which implies that $X$ is a rational homogeneous space by [Tim11, Proposition 10.1]. ∎
If the map $\Phi_{X}^{G}$ is generically finite, then $\mathcal{M}_{X}^{G}$ is the closure of a co-adjoint orbit in $\mathfrak{g}^{*}$, which carries a symplectic structure on its open orbit. In general the variety $\mathcal{M}_{X}^{G}$ is non-normal. It is then natural to pose the following:
Question 2.14.
Let $G$ be a linear algebraic group and let $X$ be a smooth projective $G$-variety. If the moment map $\Phi_{X}^{G}$ is generically finite, when is the normalization of $\mathcal{M}_{X}^{G}$ a symplectic variety?
2.5. Varieties of Fano type
Recall that a $\mathbb{Q}$-divisor $\Delta$ on a normal projective variety is called big if there exists an ample $\mathbb{Q}$-divisor $A$ and an effective $\mathbb{Q}$-divisor $N$ such that $\Delta\sim_{\mathbb{Q}}A+N$. A birational map $g\colon X\dashrightarrow X^{\prime}$ is called a birational contraction if $g^{-1}$ does not contract any divisor.
Lemma 2.15 ([PS09, Lemma-Definition 2.6 and Lemma 2.8]).
Let $X$ be a normal projective variety.
  1. (2.15.1)
    The variety $X$ is of Fano type if and only if there exists a big effective $\mathbb{Q}$-divisor $\Delta$ such that $(X,\Delta)$ is klt and $K_{X}+\Delta\sim_{\mathbb{Q}}0$.
  2. (2.15.2)
    If $X$ is of Fano type and $g\colon X\dashrightarrow X^{\prime}$ is a birational contraction to a normal projective variety, then so is $X^{\prime}$.
  3. (2.15.3)
    If $X$ is of Fano type and $g:X\to X^{\prime}$ is a projective contraction to a normal projective variety, then so is $X^{\prime}$.
Proof.
By [PS09], it remains to prove (2.15.2). Let $\Delta$ be an effective $\mathbb{Q}$-divisor such that $(X,\Delta)$ is klt and $K_{X}+\Delta\sim_{\mathbb{Q}}0$. Set $\Delta^{\prime}\coloneqq g_{*}\Delta$. Then $K_{X^{\prime}}+\Delta^{\prime}\sim_{\mathbb{Q}}0$. Let $\mu\colon W\rightarrow X$ be a log resolution of both $X$ and $g$. Let $\nu\colon W\rightarrow X^{\prime}$ be the induced birational morphism. Then the Negativity Lemma implies
$\nu^{*}(K_{X^{\prime}}+\Delta^{\prime})=\mu^{*}(K_{X}+\Delta).$
In particular, since $(X,\Delta)$ is klt and every $\nu$-exceptional divisor is also $\mu$-exceptional, the pair $(X^{\prime},\Delta^{\prime})$ is also klt and hence $X^{\prime}$ is of Fano type. ∎
We also need the following characterization of varieties of Fano type via their anti-canonical models.
Theorem 2.16 ([CG13, Theorem 1.1]).
Let $X$ be a $\mathbb{Q}$-Gorenstein normal projective variety. Then $X$ is of Fano type if and only if $-K_{X}$ is big, $R(-K_{X})$ is finitely generated and $\operatorname{Proj}R(-K_{X})$ has klt singularities.
2.6. Cox rings and Mori dream spaces
Let $X$ be a $\mathbb{Q}$-factorial normal projective variety. Choose a subgroup $\Gamma$ of Weil divisors on $X$ such that $\Gamma_{\mathbb{Q}}\rightarrow\operatorname{Cl}(X)_{\mathbb{Q}}$ is an isomorphism. Then we define the Cox ring of $X$ to be
$\operatorname{Cox}(X,\Gamma)=\bigoplus_{D\in\Gamma}H^{0}(X,\scr{O}_{X}(D))$
The definition depends on a choice of the group $\Gamma$ and if $\operatorname{Cl}(X)$ is free, then we will take $\Gamma$ such that $\Gamma\rightarrow\operatorname{Cl}(X)$ is an isomorphism and simply write $\operatorname{Cox}(X)$ for it. This is the case, for instance, if $X$ is smooth and $H^{1}(X,\scr{O}_{X})=0$. We refer the reader to [ADHL15] or [HS10, § 2] for more details about Cox rings in the general case.
Recall that a $\mathbb{Q}$-factorial normal projective variety is called a Mori dream space if its Cox ring is finitely generated [HK00]. According to [BCHM10, Corollary 1.3.2], a $\mathbb{Q}$-factorial normal projective variety of Fano type is a Mori dream space. Moreover, if $X$ is a Mori dream space and $D$ is a divisor on $X$, then there exists a Mori program for $D$; that is, a sequence of birational contractions
$X\coloneq X_{0}\dashrightarrow X_{1}\dashrightarrow\dots\dashrightarrow X_{m-1}\dashrightarrow X_{m}$
with $D_{i}$ the strict transform of $D$ on $X_{i}$ such that
  • the $X_{i}$’s are $\mathbb{Q}$-factorial normal projective varieties, and
  • the birational map $X_{i}\dashrightarrow X_{i+1}$ is either a $D_{i}$-negative divisorial contraction or a $D_{i}$-negative flip, and
  • either $D_{m}$ is nef, or $X_{m}$ has a $D_{m}$-negative elementary contraction of fiber type.
We end this section with the following result, which characterizes varieties of Fano type via their Cox ring.
Theorem 2.17 ([GOST15, Theorem 1.1]).
Let $X$ be a $\mathbb{Q}$-factorial normal projective variety. Then $X$ is of Fano type if and only if $\operatorname{Cox}(X,\Gamma)$ is finitely generated and $\operatorname{Spec}\operatorname{Cox}(X,\Gamma)$ has only klt singularities.
3. Symplectic orbifold cones
In this section we reformulate and slightly generalize the construction of symplectic orbifolds cones to the context of quasi-contact projective varieties. Let $X$ be a normal projective variety and let $L$ be a $\mathbb{Q}$-Cartier divisor on $X$. Denote by $p\colon\mathcal{L}\rightarrow X$ the natural projection (see Example 2.2) and denote by $\mathcal{L}^{\times}_{\operatorname{reg}}$ the intersection ×$\mathcal{L}^{\times}\cap\mathcal{L}_{\operatorname{reg}}$. Then $\mathcal{L}_{\operatorname{reg}}^{\times}\rightarrow X_{\operatorname{reg}}$ is a $\mathbb{C}^{\times}$-bundle equipped with a natural symplectic form induced by the contact form $\theta$ [KPSW00, § 2.2]. The dominant affinization morphism $\varphi_{L}\colon\mathcal{L}\longrightarrow\mathcal{C}_{L}$ sends the zero section $\textbf{0}_{\mathcal{L}}$ to $0\in\mathcal{C}_{L}$. This leads us to ask the following question, which is a generalization of Question 1.4.
Question 3.1.
Let $(X,L)$ be a quasi-contact normal projective variety with $L$ being a big $\mathbb{Q}$-Cartier divisor. When is $\mathcal{C}_{L}$ a symplectic variety endowed with the symplectic form induced from that on $\mathcal{L}^{\times}_{\operatorname{reg}}$?
Remark 3.2.
By [KPSW00, Theorem 1.1] and [Dem02, Corollary 4], a contact projective manifold $(X,L)$ is either a Fano manifold with Picard number one or isomorphic to $(\mathbb{P}T_{Y},\scr{O}_{\mathbb{P}T_{Y}}(1))$ for some projective manifold $Y$. So Question 1.4 can be viewed as an important and interesting special case of Question 3.1.
To answer Question 3.1, we will use the following criterion.
Proposition 3.3 ([Nam01, Theorem 6]).
Let $X$ be a normal variety. Then $X$ is a symplectic variety if and only if $X$ has only rational Gorenstein singularities and there exists a symplectic form over an open subset $U\subset X_{\operatorname{reg}}$ with $\operatorname{codim}(X\setminus U)\geq 2$
Proposition 3.4.
Let $(X,L)$ be an $n$-dimensional quasi-contact normal projective variety with $L$ being a big $\mathbb{Q}$-Cartier divisor. Then $\mathcal{C}_{L}$ is symplectic if and only if $X$ is of Fano type.
Proof.
First we assume that $X$ is of Fano type. Let $\mu\colon\widetilde{X}\rightarrow X$ be a small $\mathbb{Q}$-factorization of $X$ [BCHM10, Corollary 1.4.4]; that is, $\mu$ is a small birational contraction and $\widetilde{X}$ is $\mathbb{Q}$-factorial. Set $\widetilde{L}\coloneqq\mu^{*}L$. Then the natural pull-back
$\mu^{*}\colon H^{0}(X,\scr{O}_{X}(mL))\rightarrow H^{0}(\widetilde{X},\scr{O}_{\widetilde{X}}(m\widetilde{L}))$
is an isomorphism for any $m\in\mathbb{Z}$ and hence it yields an isomorphism $\mu^{*}\colon R(L)\rightarrow R(\widetilde{L})$ of graded $\mathbb{C}$-algebras, which satisfies the following commutative diagram
[…]
Thus, after replacing $(X,L)$ by $(\widetilde{X},\widetilde{L})$, we may assume that $X$ is $\mathbb{Q}$-factorial and so $X$ is a Mori dream space by [BCHM10, Corollary 1.3.2].
As $L$ is big, then so is $-K_{X}$. In particular, running a $-K_{X}$-MMP $g\colon X\dashrightarrow X^{\prime}$ yields a $\mathbb{Q}$-factorial normal projective variety $X^{\prime}$ with $-K_{X^{\prime}}$ big and nef. Then Lemma 2.15 says that $X^{\prime}$ is again of Fano type and hence the Basepoint Free Theorem implies that $-K_{X}$ is actually semi-ample. Denote by $X^{\prime\prime}=\operatorname{Proj}R(-K_{X^{\prime}})$ and by $h\colon X^{\prime}\rightarrow X^{\prime\prime}$ the crepant contraction. Denote $g_{*}L$ by $L^{\prime}$ and $h_{*}L^{\prime}$ by $L^{\prime\prime}$. By the Negativity Lemma, we have
$g^{*}h^{*}L^{\prime\prime}=g^{*}L^{\prime}=L-\Delta,$
where $\Delta$ is an effective $g$-exceptional divisor. For any positive integer $m$, the push-forward yields an inclusion
$\psi_{m}\colon H^{0}(X,\scr{O}_{X}(mL))\longrightarrow H^{0}(X^{\prime\prime},\scr{O}_{X^{\prime\prime}}(mL^{\prime\prime})).$
We claim that $\psi_{m}$ is actually surjective. Indeed, let $f\in K(X)=K(X^{\prime\prime})$ be a rational function such that $f\in H^{0}(\scr{O}_{X^{\prime\prime}},\scr{O}_{X^{\prime\prime}}(mL^{\prime\prime}))$, i.e., $\textup{div}(f)+mL^{\prime\prime}\geq 0$. Then we have
$\textup{div}(f)+mL=\textup{div}(f)+mg^{*}h^{*}L^{\prime\prime}+m\Delta\geq g^{*}h^{*}(\textup{div}(f)+mL^{\prime\prime})\geq 0,$
which implies that $f\in H^{0}(X,\scr{O}_{X}(mL))$. Hence, the inclusion $R(L)\rightarrow R(L^{\prime\prime})$ is actually an isomorphism of graded $\mathbb{C}$-algebras.
Since $X^{\prime}$ is of Fano type and $\mathbb{Q}$-factorial, it has only klt singularities, which then implies that $X^{\prime\prime}$ is klt as $h$ is crepant. Note that $X^{\prime\prime}$ is also of Fano type by Lemma 2.15 and the quasi-contact structure on $X$ descends to $X^{\prime\prime}$ as $h\circ g$ is a birational contraction. Hence, after replacing $(X,L)$ by $(X^{\prime\prime},L^{\prime\prime})$, we may assume furthermore that $L$ is ample. In particular, the map $\varphi_{L}\colon\mathcal{L}\rightarrow\mathcal{C}_{L}$ is a birational contraction and $\mathcal{C}_{L}$ is klt by Lemma 2.9. So the symplectic structure over $\mathcal{L}_{\operatorname{reg}}^{\times}$ descends to a symplectic structure over an open subset $U\subset\mathcal{C}_{L}$ with $\operatorname{codim}(\mathcal{C}_{L}\setminus U)\geq 2$. This implies in particular that $K_{\mathcal{C}_{L}}\sim 0$ is Cartier and hence $\mathcal{C}_{L}$ has canonical singularities. Then we conclude by Proposition 3.3
Next we assume that $\mathcal{C}_{L}$ is symplectic. Note that $\mathcal{C}_{L}$ is actually the orbifold cone associated to $(\mathbb{P}(\mathcal{C}_{L}),L^{\prime})$, where $L^{\prime}=(\widebar{\varphi}_{L})_{*}L$ and $\widebar{\varphi}_{L}\colon X\dashrightarrow\mathbb{P}(\mathcal{C}_{L})$ is the projectivization of $\varphi_{L}$. Then Lemma 2.9 implies that $\operatorname{Proj}R(-K_{X})\cong\mathbb{P}(\mathcal{C}_{L})$ is klt and hence $X$ is of Fano type by Theorem 2.16. ∎
The following claim follows from the proof.
Corollary 3.5.
Let $(X,L)$ be a quasi-contact normal projective variety with $L$ being a big $\mathbb{Q}$-Cartier divisor. If $X$ is of Fano type, then $\overline{\mathcal{C}_{L}\setminus\varphi_{L}(\mathcal{L})}^{\textup{Zar}}$ has codimension at least two in $\mathcal{C}_{L}$.
Now we are ready to finish the proof of Theorem 1.6.
Proof of Theorem 1.6.
Let $L$ be a Weil divisor such that $\scr{O}_{\mathbb{P}T_{X}}(L)\cong\scr{O}_{\mathbb{P}T_{X}}(1)$. Then there exists also an isomorphism of symplectic varieties ×$\mathcal{L}^{\times}\rightarrow T^{*}X\setminus\textbf{0}_{X}$, where $\textbf{0}_{X}$ is the zero section of $T^{*}X\rightarrow X$. Then we have the following commutative diagram:
[…]
The first statement in Theorem 1.6 is proven in Corollary 2.5. The second statement follows directly by applying Proposition 3.4 to $(\mathbb{P}T_{X},L)$, and then apply Lemma 2.15 to the projection $\mathbb{P}T_{X}\rightarrow X$. ∎
Remark 3.6.
According to the proof of Proposition 3.4, one can easily see that the conical symplectic variety $\mathcal{Z}_{X}=\mathcal{C}_{L}$ is actually an orbifold cone in the sense of Definition 2.8. This justifies our Definition 1.3.
4. Symplectic singularities from toric varieties
4.1. The Cox ring of varieties with torus action
We briefly recall the results obtained in [HS10] about the Cox ring of normal complete varieties with an algebraic torus action. Let $X$ be a $\mathbb{Q}$-factorial normal projective variety such that $\operatorname{Cl}(X)$ is finitely generated and free. Assume that $X$ carries an effective algebraic torus action $\mathbb{T}^{d}\times X\rightarrow X$. For a point $x\in X$, denote by $\mathbb{T}^{d}_{x}\subset\mathbb{T}^{d}$ the isotropy group at $x$ and consider the non-empty $\mathbb{T}^{d}$-invariant open subset
$X^{\circ}\coloneqq\{x\in X\,|\,\dim\mathbb{T}^{d}_{x}=0\}\subset X.$
By [Sum74, Corollary 2], there exists a geometric quotient $q\colon X^{\circ}\rightarrow X^{\circ}/\mathbb{T}^{d}$ with an irreducible normal but possibly non-separated orbit space $X^{\circ}/\mathbb{T}^{d}$ and then it follows from [HS10, Proposition 3.5] that there exists a seperation $f\colon X^{\circ}/\mathbb{T}^{d}\dashrightarrow Y$ to a variety $Y$, i.e., a rational map which is a local isomorphism in codimension one. Clearly both $Y$ and $f$ are unique up to isomorphisms in codimension one. Denote by $D_{1},\dots,D_{s}$ those $\mathbb{T}^{d}$-invariant prime divisors which have a finite generic isotropy group of order $>1$, and denote by $E_{1},\dots,E_{m}$ the $\mathbb{T}^{d}$-invariant prime divisors supported in $X\setminus X^{\circ}$. Then after suitably shrinking, we may assume that there exist prime divisors $C_{0},\dots,C_{r}$ on $Y$ such that each inverse image $f^{-1}(C_{j})$ is a disjoint union of prime divisors $C_{ij}$, where $1\leq j\leq n_{i}$, the map $f$ is an isomorphism over $Y\setminus(C_{0}\cup\dots\cup C_{r})$ and all the $D_{j}$’s occur among the $D_{ij}\coloneqq q^{-1}(C_{ij})$. Denote by $l_{ij}\in\mathbb{Z}$ the order of the generic isotropy group of $D_{ij}$. For a prime divisor $D$, we denote by $1_{D}\in H^{0}(X,\scr{O}_{X}(D))$ the canonical section associated to $D$.
Theorem 4.1 ([HS10, Theorem 1.2]).
There is a graded injection $\operatorname{Cox}(Y)\rightarrow\operatorname{Cox}(X)$ of Cox rings and the assignments $S_{k}\mapsto 1_{E_{k}}$ and $T_{ij}\mapsto 1_{D_{ij}}$ induce an isomorphism of $\operatorname{Cl}(X)$-graded rings
$\operatorname{Cox}(X)\cong\operatorname{Cox}(Y)[S_{1},\dots,S_{m},T_{ij};0\leq i\leq r,1\leq j\leq n_{i}]/\langle T_{i}^{l_{i}}-1_{C_{i}};0\leq i\leq r\rangle,$
where $T_{i}^{l_{i}}\coloneqq T_{i1}^{l_{i1}}\dots T_{in_{i}}^{l_{in_{i}}}$ and the $\operatorname{Cl}(X)$-grading on the right-hand side is defined by associating to $S_{k}$ the class of $E_{k}$ and to $T_{ij}$ the class of $D_{ij}$.
Note that a priori the separation $Y$ may be non-projective and $\operatorname{Cl}(Y)$ may be non-free. For the definition of Cox rings in this general case, we refer the reader to [HS10, § 2]. However, we will only use Theorem 4.1 in the case where both $X$ and $Y$ are projective manifolds with free divisor class group.
4.2. Toric hyperKähler varieties
We identify the cotangent bundle $T^{*}\mathbb{C}^{N}$ with the affine space $\mathbb{C}^{2N}$ with coordinates $(z_{1},\dots,z_{N}$, $w_{1},\dots,w_{N})$. The natural symplectic form $\omega$ on $\mathbb{C}^{2N}$ can be written as
$\omega\coloneqq\sum_{1\leq i\leq N}dz_{i}\wedge dw_{i}.$
Let $A=[\textbf{a}_{1},\dots,\textbf{a}_{N}]$ be an $(N-n)\times N$-integer matrix. Then $A$ determines a $\mathbb{T}^{N-n}$-action on $\mathbb{C}^{N}$ and hence a $\mathbb{T}^{N-n}$-action on $T^{*}\mathbb{C}^{N}=\mathbb{C}^{2N}$. More precisely, write $\textbf{t}\coloneqq(t_{1},\dots,t_{N-n})\in\mathbb{T}^{N-n}=(\mathbb{C}^{\times})^{N-n}$. Then the element t acts on $\mathbb{C}^{2N}$ as follows:
$\textbf{t}\cdot(z_{1},\dots,z_{N},w_{1},\dots,w_{N})\coloneqq(\textbf{t}^{\textbf{a}_{1}}z_{1},\dots,\textbf{t}^{\textbf{a}_{N}}z_{N},\textbf{t}^{-\textbf{a}_{1}}w_{1},\dots,\textbf{t}^{-\textbf{a}_{N}}w_{N}),$
where $\textbf{t}^{\textbf{a}_{i}}\coloneqq t_{1}^{a_{1,i}}\cdots t_{N-n}^{a_{N-n,i}}$ with $\textbf{a}_{i}=(a_{j,i})_{1\leq j\leq N-n}$. Then the moment map of the $\mathbb{T}^{N-n}$-action on $T^{*}\mathbb{C}^{N}$ is given by
$\Phi\colon T^{*}\mathbb{C}^{N}=\mathbb{C}^{2N}\longrightarrow\mathfrak{t}_{N-n}^{*}=\mathbb{C}^{N-n},\quad(z_{1},\dots,z_{N},w_{1},\dots,w_{N})\longmapsto\sum_{1\leq i\leq N}\textbf{a}_{i}z_{i}w_{i}$
From now on we assume that
[…]
  • there exists an $N\times n$-integer matrix $B$ such that no row of $B$ is zero and that the following sequence is exact:
    $0\longrightarrow\mathbb{Z}^{n}\stackrel{{\scriptstyle B}}{{\longrightarrow}}\mathbb{Z}^{N}\stackrel{{\scriptstyle A}}{{\longrightarrow}}\mathbb{Z}^{N-n}\longrightarrow 0$
Let $\theta$ be a rational character of $\mathbb{T}^{N-n}$ and $\xi\in\mathfrak{t}_{N-n}^{*}$. Then we define
$\mathcal{Y}(A,\theta,\xi)\coloneqq\Phi^{-1}(\xi)^{\theta}/\!\!/\mathbb{T}^{N-n},$
which is called a toric hyperKähler variety (aka hypertoric variety in the literature). Here $\Phi^{-1}(\xi)^{\theta}$ denotes the open subset of $\theta$-semistable points in $\Phi^{-1}(\xi)$. According to [Bel23, Proposition 2.5], the variety $\mathcal{Y}(A,\theta,\xi)$ is a symplectic variety equipped the symplectic form induced by $\omega$.
4.3. Proof of Theorem 1.8
Let $X$ be an $n$-dimensional smooth projective toric variety, given by a fan $\Sigma$ in a lattice $\mathbf{N}$, and $\mathbb{T}^{n}\cong X_{o}\subset X$ the acting torus. There are two types of rays $\rho\in\Sigma(1)$: those with $-\rho\not\in\Sigma(1)$ and those with $-\rho\in\Sigma(1)$. Denote by $\Sigma_{1}$ a subset of $\Sigma(1)$ containing all the rays of the first type and one representative for every pair of the second type. Let $\mathbf{v}_{\rho}\in\mathbf{N}$ be the primitive generator of $\rho$. We define
$\Sigma_{1}^{\perp}\coloneqq\left\{\textbf{u}\coloneqq(u_{\rho})\in\mathbb{C}^{\Sigma_{1}}\,\middle|\,\sum_{\rho\in\Sigma_{1}}u_{\rho}\mathbf{v}_{\rho}=\bf{0}\right\}.$
4.3.1. Explicit description of $\operatorname{Cox}(\mathbb{P}T_{X})$
We briefly recall the following result obtained in [HS10], which gives an explicit description of $\operatorname{Cox}(\mathbb{P}T_{X})$.
Theorem 4.2 ([HS10, Theorem 5.9]).
Let $X$ be an $n$-dimensional smooth projective toric variety given by a fan $\Sigma$. Then the Cox ring of $\mathbb{P}T_{X}$ is given by
$\operatorname{Cox}(\mathbb{P}T_{X})=\left.\mathbb{C}[S_{\rho},T_{\tau};\,\rho\in\Sigma(1),\tau\in\Sigma_{1}\,]\middle/\left\langle\sum_{\rho\in\Sigma_{1}}u_{\rho}S^{\rho}T_{\rho};\,\textbf{u}=(u_{\rho})\in\Sigma_{1}^{\perp}\right\rangle\right.$
where
$S^{\rho}\coloneqq\begin{cases}S_{\rho}S_{-\rho},&-\rho\in\Sigma(1);\\ S_{\rho},&\textup{else}.\end{cases}$
To simplify the notation, we set $\mathcal{X}\coloneqq\mathbb{P}T_{X}$. There exists an open subset $\mathcal{X}^{\circ}$ of $\mathcal{X}$ with $\operatorname{codim}(\mathcal{X}\setminus\mathcal{X}^{\circ})\geq 2$ such that $\mathbb{T}^{n}$ acts on $\mathcal{X}^{\circ}$ with trivial isotropy subgroup and $\mathbb{P}(\mathfrak{t}_{n})$ is a separation of $\mathcal{X}^{\circ}/\mathbb{T}^{n}$ such that the projectivized moment map
$\widebar{\Phi}\colon\mathcal{X}=\mathbb{P}T_{X}\dashrightarrow\mathbb{P}(\mathfrak{t}_{n})$
coincides with the composition
$\mathcal{X}\stackrel{{\scriptstyle q}}{{\dashrightarrow}}\mathcal{X}^{\circ}/\mathbb{T}^{n}\stackrel{{\scriptstyle f}}{{\dashrightarrow}}\mathbb{P}(\mathfrak{t}_{n}).$
Note that $\mathbf{v}_{\rho}$ can be regarded as an element in $\mathbb{T}^{n}=\mathbf{N}\otimes_{\mathbb{Z}}\mathbb{T}$ and therefore an element in $\mathfrak{t}_{n}$. Denote by $C_{\rho}\subset\mathbb{P}(\mathfrak{t}_{n})$ the hyperplane defined by the natural quotient $\mathfrak{t}_{n}\rightarrow\mathfrak{t}_{n}/\mathbb{C}\mathbf{v}_{\rho}$. Then the Cox ring of the projective space $\mathbb{P}(\mathfrak{t}_{n})$ can be naturally identified to
$\left.\mathbb{C}[U_{\rho};\,\rho\in\Sigma_{1}]\middle/\left\langle\sum_{\rho\in\Sigma_{1}}u_{\rho}U_{\rho};\,\textbf{u}=(u_{\rho})\in\Sigma_{1}^{\perp}\right\rangle\right.,$
where $U_{\rho}$ corresponds to the canonical section $1_{C_{\rho}}$. The pull-back $\widebar{\Phi}^{*}C_{\rho}$ contains a unique $\widebar{\pi}$-horizontal prime divisor, denoted by $H_{\rho}$. Then we have
$\widebar{\Phi}^{*}C_{\rho}=\begin{cases}H_{\rho}+\widebar{\pi}^{*}D_{\rho}+\widebar{\pi}^{*}D_{-\rho},&-\rho\in\Sigma(1);\\ H_{\rho}+\widebar{\pi}^{*}D_{\rho},&\textup{else};\end{cases}$
where $D_{\rho}$ is the toric divisor on $X$ corresponding to $\rho$. Then Theorem 4.2 follows from Theorem 4.1 by identifying $S_{\rho}$ (resp. $T_{\tau}$) to the canonical section $1_{\widebar{\pi}^{*}D_{\rho}}$ (resp. $1_{H_{\tau}}$).
4.3.2. Characterize elements in $S(X)$
Now we aim to characterize the elements of $\operatorname{Cox}(\mathbb{P}T_{X})$ contained in the subring $S(X)$. To this end, we shall interpret $\operatorname{Cox}(\mathbb{P}T_{X})$ as the following ring:
$R$ $\coloneqq\left.\mathbb{C}[S_{\rho},T_{\rho};\,\rho\in\Sigma(1)\,]\middle/\left\langle\sum_{\rho\in\Sigma_{1}}u_{\rho}S^{\rho}T_{\rho},(u_{\rho})\in\Sigma_{1}^{\perp};T_{\rho}+T_{-\rho},\rho\in\Sigma(1)\setminus\Sigma_{1}\right\rangle\right.$
$\cong\operatorname{Cox}(\mathbb{P}T_{X})$
Let $\mathbf{M}$ be the lattice of characters of $\mathbb{T}^{n}$. Set $N=|\Sigma(1)|$. Then we have the following short exact sequence
$0\longrightarrow\mathbf{M}\stackrel{{\scriptstyle B}}{{\longrightarrow}}\mathbb{Z}^{\Sigma(1)}=\mathbb{Z}^{N}\stackrel{{\scriptstyle A}}{{\longrightarrow}}\operatorname{Pic}(X)\longrightarrow 0$ (4.2.1)
Denote by $\ker(A)\subset\mathbb{Z}^{\Sigma(1)}$ the kernel of the map $A$. Consider the following surjective map
$\mathbb{C}[S_{\rho},T_{\rho};\rho\in\Sigma(1)]\longrightarrow R.$
Given $I^{\prime}=(i^{\prime}_{\rho}),I=(i_{\rho})\in\mathbb{Z}_{\geq 0}^{\Sigma(1)}$, we denote by $S^{I^{\prime}}T^{I}$ the monomial
$\prod_{\rho\in\Sigma(1)}S_{\rho}^{i^{\prime}_{\rho}}T_{\rho}^{i_{\rho}}.$
By abuse of notation, we shall also use $S^{I^{\prime}}T^{I}$ to denote its image of $R$.
Lemma 4.3.
$S^{I^{\prime}}T^{I}\in S(X)$ if and only if $I^{\prime}-I-\prescript{c}{}{I}\in\ker(A)$, where $\prescript{c}{}{I}=(\prescript{c}{}{i_{\rho}})$ is defined as follows:
$\prescript{c}{}{i_{\rho}}\coloneqq\begin{cases}i_{-\rho}&\textup{if}\quad-\rho\in\Sigma(1);\\ 0&\textup{else}.\end{cases}$
Proof.
By Theorem 4.2 and its proof, the element $S^{I^{\prime}}T^{I}$ can be viewed as the canonical section of $\scr{O}_{\mathbb{P}T_{X}}(H_{I}+\widebar{\pi}^{*}D_{I^{\prime}})$ with
$H_{I}\coloneqq\sum_{\rho\in\Sigma(1)}i_{\rho}H_{\rho},\quad D_{I^{\prime}}\coloneqq\sum_{\rho\in\Sigma(1)}i^{\prime}_{\rho}D_{\rho}$
where $H_{\rho}=H_{-\rho}$ if $\rho\in\Sigma(1)\setminus\Sigma_{1}$. As $\scr{O}_{\mathbb{P}T_{X}}(1)\sim H_{\rho}+\widebar{\pi}^{*}(D_{\rho}+D_{-\rho})$ for any $\rho\in\Sigma(1)$, where $D_{-\rho}\coloneqq 0$ if $-\rho\not\in\Sigma(1)$, the monomial $S^{I^{\prime}}T^{I}$ corresponds to a section of
$\operatorname{Sym}^{m}T_{X}\otimes\scr{O}_{X}(D_{I^{\prime}}-D_{I}-D_{\prescript{c}{}{I}}),$
where $m=\sum i_{\rho}$. So $S^{I^{\prime}}T^{I}$ is contained in $S(X)$ if and only if $D_{I^{\prime}}\sim D_{I}+D_{\prescript{c}{}{I}}$, which is equivalent to $I^{\prime}-I-\prescript{c}{}{I}\in\ker(A)$ by (4.2.1). ∎
Before giving the proof of Theorem 1.8, we discuss a concrete example to illustrate the idea.
Example 4.4.
Let $X\rightarrow\mathbb{P}^{n}$ be the blow-up of $\mathbb{P}^{n}$ at the point $p\coloneqq[0:\dots:0:1]$. Then $X$ is toric and its fan has ray generators given by $\textbf{v}_{i}=\textbf{e}_{i}$ for $1\leq i\leq n$, $\mathbf{v}_{n+1}=-\textbf{e}_{1}-\dots-\textbf{e}_{n}$ and $\textbf{v}_{n+2}=-\textbf{v}_{n+1}$. We choose $\Sigma_{1}$ to be $\{\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}\}$. Then $\Sigma_{1}^{\perp}=\mathbb{C}\mathbf{u}$, where $\mathbf{u}=(1,\dots,1)$, and Theorem 4.2 implies
$\operatorname{Cox}(\mathbb{P}T_{X})\cong\left.\mathbb{C}[x_{1},\dots,x_{n+2},y_{1},\dots,y_{n+1}]\middle/\left\langle\sum_{i=1}^{n}x_{i}y_{i}+x_{n+1}x_{n+2}y_{n+1}\right\rangle\right.$
Now we add another variable $y_{n+2}$ and write $\operatorname{Cox}(\mathbb{P}T_{X})$ as
$R\coloneqq\left.\mathbb{C}[x_{1},\dots,x_{n+2},y_{1},\dots,y_{n+2}]\middle/\left\langle\sum_{i=1}^{n}x_{i}y_{i}+x_{n+1}x_{n+2}y_{n+1},y_{n+1}+y_{n+2}\right\rangle\right.$
Note that the matrix $A$ associated to $X$ is
$\begin{pmatrix}1&\dots&1&1&0\\ -1&\dots&-1&0&1\end{pmatrix}.$
So the toric hyperKähler variety $\mathcal{Y}(A,0,0)$ associated to $A$ is the spectrum of the $\mathbb{T}^{2}$-invariant subring of the following ring $R^{\prime}$:
$R^{\prime}\coloneqq\left.\mathbb{C}[x^{\prime}_{1},\dots,x^{\prime}_{n+2},y^{\prime}_{1},\dots,y^{\prime}_{n+2}]\middle/\left\langle\sum_{i=1}^{n+1}x^{\prime}_{i}y^{\prime}_{i},\ -\sum_{i=1}^{n}x^{\prime}_{i}y^{\prime}_{i}+x^{\prime}_{n+2}y^{\prime}_{n+2}\right\rangle\right..$
We also note that $R^{\prime}$ can also be naturally identified to the following ring
$\widetilde{R}^{\prime}\coloneqq\left.\mathbb{C}[x^{\prime}_{1},\dots,x^{\prime}_{n+2},y^{\prime}_{1},\dots,y^{\prime}_{n+2}]\middle/\left\langle\sum_{i=1}^{n+1}x^{\prime}_{i}y^{\prime}_{i},\ x^{\prime}_{n+1}y^{\prime}_{n+1}+x^{\prime}_{n+2}y^{\prime}_{n+2}\right\rangle\right..$
Let $\widetilde{R}$ be the ring defined as follows:
$\left.\mathbb{C}[x_{1},\dots,x_{n+2},y_{1},\dots,y_{n+2}]\middle/\left\langle\sum_{i=1}^{n}x_{i}y_{i}+x_{n+1}x_{n+2}y_{n+1},\ x_{n+1}x_{n+2}(y_{n+1}+y_{n+2})\right\rangle\right.$
Then we get a natural commutative diagram
[…]
where $\widetilde{\phi}$ is defined as follows:
$\begin{cases}x^{\prime}_{i}\mapsto x_{i}&1\leq i\leq n+2;\\ y^{\prime}_{i}\mapsto y_{i}&1\leq i\leq n;\\ y^{\prime}_{n+1}\mapsto x_{n+2}y_{n+1}&\\ y^{\prime}_{n+2}\mapsto x_{n+1}y_{n+2}.&\end{cases}$
Then $\widebar{\phi}^{*}\colon\operatorname{Spec}(R)\rightarrow\operatorname{Spec}(R^{\prime})$ is birational and the $\mathbb{T}^{2}$-action on $R^{\prime}$ can be naturally extended to $R$ with
$\operatorname{wt}(y_{n+1})=(-1,-1)\quad\textup{and}\quad\operatorname{wt}(y_{n+2})=(-1,-1).$
Let $x^{I}y^{J}\coloneqq x_{1}^{i_{1}}\dots x_{n+2}^{i_{n+2}}y_{1}^{j_{1}}\dots y_{n+2}^{j_{n+2}}$ be a monomial. By Lemma 4.3, one can derive that $x^{I}y^{J}\in S(X)$ if and only if the following equalities hold:
$\begin{dcases}\sum_{k=1}^{n+1}(i_{k}-j_{k})-j_{n+2}=0\\ \sum_{k=1}^{n}(j_{k}-i_{k})+i_{n+2}-j_{n+1}-j_{n+2}=0.\end{dcases}$
This yields
$i_{n+1}+i_{n+2}=2(j_{n+1}+j_{n+2}).$
As $y_{n+2}=-y_{n+1}$ in $R$, after suitable replacement, we may assume that $i_{n+1}\geq j_{n+2}$ and $i_{n+2}\geq j_{n+1}$; that is, $x^{I}y^{J}$ is contained in the image $\widebar{\phi}$. This implies
$S(X)\cong R^{\mathbb{T}^{2}}=(R^{\prime})^{\mathbb{T}^{2}}.$
4.3.3. Auxiliary lemmas
Fix a basis of $\operatorname{Pic}(X)$ such that $\operatorname{Pic}(X)\cong\mathbb{Z}^{N-n}$. Then we can write $A=[\textbf{a}_{\rho};\rho\in\Sigma(1)]$ and $\prescript{t}{}{A}=[\widebar{\mathbf{a}}_{i};1\leq i\leq N-n]$. Note that the linear map $B$ is given as
$\mathbf{M}\longrightarrow\mathbb{Z}^{\Sigma(1)},\quad\mathbf{m}\longmapsto(\langle\mathbf{v}_{\rho},\mathbf{m}\rangle)_{\rho\in\Sigma(1)},$
so the exactness of (4.2.1) implies that the vector space
$\Sigma(1)^{\perp}\coloneqq\left\{\mathbf{c}=(c_{\rho})\in\mathbb{C}^{\Sigma(1)}\,\middle|\,\sum_{\rho\in\Sigma(1)}c_{\rho}\mathbf{v}_{\rho}=0\right\}$
is generated by the $\widebar{\mathbf{a}}_{i}$’s $(1\leq i\leq N-n)$. Moreover, for any $\rho\in\Sigma(1)\setminus\Sigma_{1}$, we define an element
$\prescript{\rho}{}{\mathbf{c}}\coloneqq(\prescript{\rho}{}{c}_{\rho^{\prime}})\in\mathbb{C}^{\Sigma(1)}$
with
$\prescript{\rho}{}{c}_{\rho^{\prime}}\coloneqq\begin{cases}1&\text{if}\quad\rho^{\prime}=\rho\quad\text{or}\quad-\rho;\\ 0&\text{else}.\end{cases}$
Then the exactness of (4.2.1) implies that $\prescript{\rho}{}{\mathbf{c}}$ is contained in $\Sigma(1)^{\perp}$.
Lemma 4.5.
The space $\Sigma(1)^{\perp}$ is generated by $\Sigma_{1}^{\perp}$ and the $\prescript{\rho}{}{\mathbf{c}}$’s, where $\Sigma_{1}^{\perp}\hookrightarrow\Sigma(1)^{\perp}$ is given as
$\mathbf{u}\longmapsto(\mathbf{u},0,\dots,0).$
Proof.
This follows from the fact that $\mathbf{v}_{\rho}+\mathbf{v}_{-\rho}=0$ if $\rho\in\Sigma(1)\setminus\Sigma_{1}$. ∎
Now we introduce a ring $R^{\prime}$ as follows:
$R^{\prime}\coloneqq\left.\mathbb{C}[S_{\rho},T^{\rho};\,\rho\in\Sigma(1)\,]\middle/\left\langle\sum_{\rho\in\Sigma(1)}\mathbf{a}_{\rho}S_{\rho}T^{\rho}\right\rangle\right.$
Then we have $\operatorname{Spec}R^{\prime}\cong\Phi^{-1}(0)$ and hence $R^{\prime}$ is a domain by [BK12, Lemma 4.7]. On the other hand, consider the following natural morphism
$\phi\colon\mathbb{C}[S_{\rho},T^{\rho};\rho\in\Sigma(1)]\longrightarrow\mathbb{C}[S_{\rho},T_{\rho};\rho\in\Sigma(1)]$
with
$T^{\rho}\longmapsto\begin{cases}S_{-\rho}T_{\rho}&\textup{if}\quad-\rho\in\Sigma(1);\\ T_{\rho}&\textup{else}.\end{cases}$
Lemma 4.6.
The morphism $\phi$ induces a morphism $\widebar{\phi}\colon R^{\prime}\rightarrow R$ such that the induced morphism $\widebar{\phi}^{*}\colon\operatorname{Spec}R\rightarrow\operatorname{Spec}R^{\prime}$ is birational.
Proof.
Firstly we show that $\widebar{\phi}$ is well-defined. To see this, consider the ring $\widetilde{R}$ defined as follows:
$\left.\mathbb{C}[S_{\rho},T_{\rho};\,\rho\in\Sigma(1)\,]\middle/\left\langle\sum_{\rho\in\Sigma_{1}}u_{\rho}S^{\rho}T_{\rho},(u_{\rho})\in\Sigma_{1}^{\perp};S^{\rho}(T_{\rho}+T_{-\rho}),\rho\in\Sigma(1)\setminus\Sigma_{1}\right\rangle\right.$
Then we have a natural quotient homomorphism $\widetilde{R}\rightarrow R$. On the other hand, consider the ring $\widetilde{R}^{\prime}$ defined as follows:
$\left.\mathbb{C}[S_{\rho},T^{\rho};\,\rho\in\Sigma(1)\,]\middle/\left\langle\sum_{\rho\in\Sigma_{1}}u_{\rho}S_{\rho}T^{\rho},(u_{\rho})\in\Sigma_{1}^{\perp};S_{\rho}T^{\rho}+S_{-\rho}T^{-\rho},\rho\in\Sigma(1)\setminus\Sigma_{1}\right\rangle\right.$
Then Lemma 4.5 implies that $\widetilde{R}^{\prime}$ is isomorphic to the following ring:
$\left.\mathbb{C}[S_{\rho},T^{\rho};\,\rho\in\Sigma(1)\,]\middle/\left\langle\sum_{\rho\in\Sigma(1)}c_{\rho}S_{\rho}T^{\rho};\,\mathbf{c}=(c_{\rho})\in\Sigma(1)^{\perp}\right\rangle\right.$
In particular, since $\Sigma(1)^{\perp}$ is generated by the $\widebar{\mathbf{a}}_{i}$’s $(1\leq i\leq N-n)$, the ring $\widetilde{R}^{\prime}$ is canonically identified to $R^{\prime}$ and we have the following commutative diagram
[…]
where $\widetilde{\phi}$ is the natural morphism induced by $\phi$. Hence $\widebar{\phi}$ is well-defined.
Note that $S_{\rho}\not=0$ in $R$ and and hence so is in $R^{\prime}$ for any $\rho\in\Sigma(1)$. In particular, since $R$ and $R^{\prime}$ are domains, the rational map $(\widebar{\phi}^{*})^{-1}\colon\operatorname{Spec}R^{\prime}\dashrightarrow\operatorname{Spec}R$ defined by sending $T_{\rho}$ to $T^{\rho}/S_{-\rho}$ if $\rho\in\Sigma(1)\setminus\Sigma_{1}$ shows that $\widebar{\phi}^{*}$ is indeed birational. ∎
4.3.4. Proof of Theorem 1.8
The torus $\mathbb{T}^{N-n}$ acts on $R^{\prime}$ with $\operatorname{wt}(S_{\rho})=\mathbf{a}_{\rho}$ and $\operatorname{wt}(T^{\rho})=-\mathbf{a}_{\rho}$ for any $\rho\in\Sigma(1)$. Note that $R^{\prime}$ is a subalgebra of $R$ by Lemma 4.6. Then the $\mathbb{T}^{N-n}$-action can be extended to $R$ such that $\operatorname{wt}(T_{\rho})=-\mathbf{a}_{\rho}-\mathbf{a}_{-\rho}$ for any $\rho\in\Sigma(1)\setminus\Sigma_{1}$. Then an easy computation shows that a monomial $S^{I^{\prime}}T^{I}\in R$, where $I^{\prime}=(i^{\prime}_{\rho}),I=(i_{\rho})\in\mathbb{Z}_{\geq 0}^{\Sigma(1)}$, has weight 0 if and only if $I^{\prime}-I-\prescript{c}{}{I}\in\ker(A)$ and the latter is equivalent to $S^{I^{\prime}}T^{I}\in S(X)$ by Lemma 4.3. So it is enough to show that every $\mathbb{T}^{N-n}$-invariant monomial $S^{I^{\prime}}T^{I}$ in $R$ is contained in $R^{\prime}$.
Let $S^{I^{\prime}}T^{I}$ be a $\mathbb{T}^{N-n}$-invariant element in $R$. Then $I^{\prime}-I-\prescript{c}{}{I}\in\ker(A)$, or equivalently $I^{\prime}-I-\prescript{c}{}{I}\in\textup{Im}(B)$. In particular, as $\mathbf{v}_{\rho}=-\mathbf{v}_{-\rho}$ for any $\rho\in\Sigma(1)\setminus\Sigma_{1}$, we get
$i^{\prime}_{\rho}-i_{\rho}-i_{-\rho}=-i^{\prime}_{-\rho}+i_{-\rho}+i_{-\rho}$
for any $\rho$ such that $-\rho\in\Sigma(1)$. In particular, as $T_{\rho}=-T_{-\rho}$ in $R$, we may assume that $i^{\prime}_{\rho}\geq i_{-\rho}$ if $-\rho\in\Sigma(1)$ and hence the monomial $S^{I^{\prime}}T^{I}$ is contained in the subring $R^{\prime}$ of $R$ generated by the $S_{\rho}$’s and the $S_{-\rho}T_{\rho}=T^{\rho}$’s. ∎
Remark 4.7.
It is noteworthy that extending Theorem 1.8 to any toric variety $X$ with $H^{0}(X,\scr{O}_{X}^{\times})=\mathbb{C}^{\times}$ and $\operatorname{Pic}(X)$ free poses no essential difficulties (see [ADHL15, Theorem 4.4.1.3]). However, given that the primary focus of this article is to present a distinctive perspective for studying $\mathcal{Z}_{X}$ or equivalently $S(X)$, we choose to concentrate on smooth projective varieties.
4.4. Example: blow-ups of projective spaces along points
Note that $\mathcal{Z}_{\mathbb{P}^{n}}$ is the minimal nilpotent orbit closure $\overline{\mathcal{O}}_{\textup{min}}\subset\mathfrak{sl}_{n+1}$ consisting of traceless matrices of rank at most 1. The map $\Phi_{\mathbb{P}^{n}}:T^{*}\mathbb{P}^{n}\to\overline{\mathcal{O}}_{\textup{min}}$ is the Springer resolution, which collapses the zero section to the origin. This map can be explicitly described as follows: Let $\mathbb{T}$ act on $\mathbb{C}^{n+1}\times(\mathbb{C}^{n+1})^{*}$ by
$\lambda\cdot(x_{1},\cdots,x_{n+1},y_{1},\cdots,y_{n+1})=(\lambda x_{1},\cdots,\lambda x_{n+1},\lambda^{-1}y_{1},\cdots,\lambda^{-1}y_{n+1}).$
This $\mathbb{T}$-action preserves the hypersurface $\mathcal{H}\coloneqq\{(\textbf{x},\textbf{y})|\sum x_{i}y_{i}=0\}$. Moreover, since $\mathbb{P}^{n}$ is a toric variety given by the fan $\Sigma_{\mathbb{P}^{n}}$, which has rays generated by $\mathbf{v}_{i}=\textbf{e}_{i}$, $1\leq i\leq n$, and $\mathbf{v}_{n+1}=-\textbf{e}_{1}-\dots-\textbf{e}_{n}$, then Theorem 4.2 implies that $\operatorname{Spec}S(\mathbb{P}^{n})\cong\mathcal{H}/\!\!/\mathbb{T}$. The Springer resolution $\Phi_{\mathbb{P}^{n}}$ is then given by
$({\bf x},{\bf y})\mapsto(x_{i}y_{j})_{1\leq i,j\leq n+1}.$
Now let $X$ be the blow-up of $\mathbb{P}^{n}$ at the toric point $[0:\dots:0:1]$. We will describe in details the symplectic singularity $\mathcal{Z}_{X}$ (see Example 4.4). The toric structure of $X$ gives a $\mathbb{T}^{2}$-action on $\mathbb{C}^{n+2}$ via
$(t_{1},t_{2})\cdot(x_{1},\cdots,x_{n+2})=(t_{1}t_{2}^{-1}x_{1},\cdots,t_{1}t_{2}^{-1}x_{n},t_{1}x_{n+1},t_{2}x_{n+2}).$
The associated moment map is then given by
$\Phi\colon\mathbb{C}^{2n+4}\to\mathbb{C}^{2},\quad(x_{i},y_{j})\mapsto\left(\sum_{i=1}^{n}x_{i}y_{i}+x_{n+1}y_{n+1},-\sum_{i=1}^{n}x_{i}y_{i}+x_{n+2}y_{n+2}\right).$
The coordinate ring of the quotient $\Phi^{-1}(0)/\!\!/\mathbb{T}^{2}$ is generated by the following monomials:
$x_{i}y_{j}\,(1\leq i,j\leq n),\quad x_{i}y_{n+1}x_{n+2}\,(1\leq i\leq n),\quad y_{i}x_{n+1}y_{n+2}\,(1\leq i\leq n).$
It follows from Theorem 1.8 that $\mathcal{Z}_{X}\cong\Phi^{-1}(0)/\!\!/\mathbb{T}^{2}=\mathcal{Y}(A,0,0)$, which is isomorphic to the following
$\left\{(z_{ij})_{1\leq i,j\leq n+1}\,\left|\,\sum_{i=1}^{n+1}z_{ii}=0,\quad{\rm rk}\begin{pmatrix}z_{1,1}&\cdots&z_{1,n}&z_{1,n+1}\\ \vdots&\vdots&\vdots&\vdots\\ z_{n,1}&\cdots&z_{n,n}&z_{n,n+1}\\ z_{n+1,1}&\cdots&z_{n+1,n}&-z_{n+1,n+1}^{2}\end{pmatrix}\leq 1\right\}\right..$
Note that when $n=2$, this example is studied in [Nag21, Theorem 4.8]. The singular locus of $\mathcal{Z}_{X}$ is given by
$S\coloneqq\left.\left\{C=\begin{pmatrix}z_{1,1}&\cdots&z_{1,n}&0\\ \vdots&\vdots&\vdots&\vdots\\ z_{n,1}&\cdots&z_{n,n}&0\\ 0&\cdots&0&0\end{pmatrix}\,\right|\,{\rm tr}(C)=0,{\rm rk}(C)\leq 1\right\}$
Note that $S$ is isomorphic to the minimal nilpotent orbit closure in $\mathfrak{sl}_{n}$. Write $S=S_{0}\cup\{0\}$ as the union of rank 1 (resp. 0) traceless matrices. Note that $S$ has codimension 2 in $\mathcal{Z}_{X}$ and the transverse singularity of $\mathcal{Z}_{X}$ along any point of $S_{0}$ is an $A_{1}$-singularity. The blow-up $X\to\mathbb{P}^{n}$ induces a natural morphism
$H^{0}(T^{*}X,\mathcal{O}_{T^{*}X})\to H^{0}(T^{*}\mathbb{P}^{n},\mathcal{O}_{T^{*}\mathbb{P}^{n}}),$
which gives a map $\nu\colon\overline{\mathcal{O}}_{\textup{min}}\to\mathcal{Z}_{X}$. In terms of matrices, this map is given by
$z\coloneqq\begin{pmatrix}z_{1,1}&\cdots&z_{1,n+1}\\ \vdots&\vdots&\vdots\\ z_{n+1,1}&\cdots&z_{n+1,n+1}\end{pmatrix}\mapsto\nu(z)\coloneqq\begin{pmatrix}z_{1,1}&\cdots&z_{1,n}&-z_{n+1,n+1}z_{1,n+1}\\ \vdots&\vdots&\vdots&\vdots\\ z_{n,1}&\cdots&z_{n,n}&-z_{n+1,n+1}z_{n,n+1}\\ z_{n+1,1}&\cdots&z_{n+1,n}&z_{n+1,n+1}\end{pmatrix}.$
The divisor $\{z_{n+1,n+1}=0\}$ in $\overline{\mathcal{O}}_{\textup{min}}$ has two components: $D_{1}$ (the last row is zero) and $D_{2}$ (the last column is zero). One can easily check that:
  1. (4.7.1)
    The map $\nu$ is an isomorphism outside $D_{1}\cup D_{2}$.
  2. (4.7.2)
    The image $\nu(D_{1})$ is the singular locus $S$ of $\mathcal{Z}_{X}$. The fibers of $\nu|_{D_{1}}\colon D_{1}\to S$ are $\mathbb{A}^{1}$ outside the origin and $\mathbb{A}^{n}$ over the origin.
  3. (4.7.3)
    The map $\nu|_{D_{2}}\colon D_{2}\to\nu(D_{2})$ is an isomorphism.
  4. (4.7.4)
    The map $\nu$ is not surjective. The closure of $\mathcal{Z}_{X}\setminus{\rm Im}(\nu)$ is a divisor $W$ consisting of matrices with the last row being identically zero. Moreover, we have $\mathcal{Z}_{X}\setminus{\rm Im}(\nu)=W\setminus S$.
As the matrix $A$ is unimodular, we have a symplectic resolution given by $\mathcal{Y}(A,\theta,0)\to\mathcal{Y}(A,0,0)$ for a general $\theta$. Write $A=[\textbf{a}_{1},\dots,\textbf{a}_{n+2}]$ into columns, namely
$\textbf{a}_{1}=\cdots=\textbf{a}_{n}=\begin{pmatrix}1\\ -1\end{pmatrix},\textbf{a}_{n+1}=\begin{pmatrix}1\\ 0\end{pmatrix},\textbf{a}_{n+2}=\begin{pmatrix}0\\ 1\end{pmatrix}.$
The $\theta$-semistable set is given by
$(\mathbb{C}^{2n+4})^{\theta-ss}=\{(\textbf{x},\textbf{y})\in\mathbb{C}^{2n+4}|\theta\in\sum_{i:x_{i}\neq 0}\mathbb{Q}_{\geq 0}\textbf{a}_{i}+\sum_{j:y_{j}\neq 0}\mathbb{Q}_{\geq 0}(-\textbf{a}_{j})\}.$
Choose $\theta=\begin{pmatrix}1\\ -1\end{pmatrix}$, then we have
$\Phi^{-1}(0)^{\theta}=\{(\textbf{x},\textbf{y})\in\mathbb{C}^{2n+4}|(x_{1},\cdots,x_{n+1})\neq 0,y_{n+2}\neq 0\}\cap\Phi^{-1}(0).$
Then $\mathcal{Y}(A,\theta,0)=\Phi^{-1}(0)^{\theta}//\mathbb{T}^{2}$ and the symplectic resolution $\beta:\mathcal{Y}(A,\theta,0)\to\mathcal{Y}(A,0,0)$ is induced by
$\Phi^{-1}(0)^{\theta}\ni(\textbf{x},\textbf{y})\mapsto\begin{pmatrix}x_{1}y_{1}&\cdots&x_{1}y_{n}&-x_{1}y_{n+1}x_{n+2}\\ \vdots&\vdots&\vdots&\vdots\\ x_{n}y_{1}&\cdots&x_{n}y_{n}&-x_{n}y_{n+1}x_{n+2}\\ x_{n+1}y_{1}y_{n+2}&\cdots&x_{n+1}y_{n}y_{n+2}&x_{n+1}y_{n+1}\end{pmatrix}.$
Now we define an injective morphism $\mathcal{H}\to\Phi^{-1}(0)^{\theta}$ as follows:
$\mathbb{C}^{2n+2}\supset\mathcal{H}\ni(\textbf{x},\textbf{y})\mapsto(x_{1},\cdots,x_{n+1},-x_{n+1}y_{n+1},y_{1},\cdots,y_{n+1},1).$
This map descends to the quotients to give an injective morphism
$\tilde{\nu}:T^{*}_{\mathbb{P}^{n}}\to\mathcal{Y}(A,\theta,0),$
which makes the following diagram commutative:
${T^{*}_{\mathbb{P}^{n}}}$${\mathcal{Y}(A,\theta,0)}$${\overline{\mathcal{O}}_{\textup{min}}}$${\mathcal{Z}_{X}\simeq\mathcal{Y}(A,0,0)}$$\scriptstyle{\widetilde{\nu}}$$\scriptstyle{\Phi_{\mathbb{P}^{n}}}$$\scriptstyle{\beta}$$\scriptstyle{\nu}$
Now we compute the central fiber of $\beta$. Assume $(\textbf{x},\textbf{y})\in\Phi^{-1}(0)^{\theta}$ which is mapped to 0 by $\tilde{\nu}$. There are two cases:
Case (i): there exists an index $1\leq i\leq n$ such that $x_{i}\neq 0$. Then we get $y_{1}=\cdots=y_{n}=0$ and $x_{n+1}y_{n+1}=x_{n+2}y_{n+2}=y_{n+1}x_{n+2}=0$. As $y_{n+2}\neq 0$, we obtain two components in this case, both are isomorphic to $\mathbb{P}^{n}$:
$E_{1}=\{x_{n+2}=0,y_{1}=\cdots=y_{n}=y_{n+1}=0\}//\mathbb{T}^{2},$
which is in fact the image of the zero-section $\mathbf{0}_{\mathbb{P}^{n}}$, and
$E_{2}=\{x_{n+1}=x_{n+2}=0,y_{1}=\cdots=y_{n}=0\}//\mathbb{T}^{2},$
which is in face the compactification of the $\mathbb{A}^{n}$-fiber of $\nu|_{D_{1}}$ over the origin.
Case (ii): $x_{1}=\cdots=x_{n}=0$. Then we have $x_{n+1}\neq 0$ (as $(x_{1},\cdots,x_{n+1})\neq 0).$ Recall that $y_{n+2}\neq 0$, we must have $y_{1}=\cdots=y_{n}=y_{n+1}=0$ and $x_{n+2}=0$. But this subset is already contained in case (i).
In conclusion, the central fiber of $\beta$ consists of two $\mathbb{P}^{n}$’s. We can obtain two more symplectic resolutions by performing Mukai flops along these two $\mathbb{P}^{n}$’s. These two symplectic resolutions correspond to different choices of $\theta$. This give three symplectic resolutions for $\mathcal{Z}_{X}$.
In general, if $X$ is the blow-up of $n+1-k$ toric fixed points of $\mathbb{P}^{n}$ (with $k\leq n+1$), then the symplectic singularity $\mathcal{Z}_{X}$ is then isomorphic to the following
$\left\{(z_{ij})_{1\leq i,j\leq n+1}\,\left|\,\sum_{i=1}^{n+1}z_{ii}=0,\quad{\rm rk}\begin{pmatrix}z_{1,1}&\cdots&\cdots&\cdots&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ z_{k,1}&\cdots&z_{k,k}&\cdots&\cdots\\ z_{k+1,1}&\cdots&\cdots&-z^{2}_{k+1,k+1}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ z_{n+1,1}&\cdots&\cdots&\cdots&-z_{n+1,n+1}^{2}\end{pmatrix}\leq 1\right\}\right..$
We leave it to the readers to describe the symplectic resolutions in this general case.
5. Symplectic singularities from horospherical varieties
5.1. Horospherical varieties and their moment maps
Let $G$ be a reductive and connected linear algebraic group and let $B$ be a Borel subgroup of $G$. A closed subgroup $H$ of $G$ is said to be horospherical if it contains the unipotent radical of a Borel subgroup of $G$ and we also say that the homogeneous space $G/H$ is horospherical. Denote by $P$ the normalizer $N_{G}(H)$ of $H$ in $G$. Then $P$ is a parabolic subgroup of $G$ such that $P/H$ is a torus. Thus $G/H$ is a torus bundle over the flag variety $W\coloneqq G/P$. The dimension $r$ of $P/H$ is called the rank of $G/H$. A normal variety $X$ with an action of $G$ is said to be a horospherical variety if $G$ has an open orbit isomorphic to $G/H$ for some horospherical subgroup $H$.
Proposition 5.1 ([Pas16, Proposition 4.1]).
Let $X$ be a horospherical $G$-variety with open orbit $X_{o}=G/H$ isomorphic to a torus fibration $X_{o}\rightarrow G/P$ with $P$ a parabolic subgroup of $G$. Then, there exists a smooth toric $P/H$-variety $Y$ and a $G$-equivariant birational morphism $f$ from the smooth horospherical $G$-variety $G\times_{P}Y$ to $X$.
Let $\mathfrak{g}$, $\mathfrak{h}$ and $\mathfrak{p}$ be the Lie algebras of $G$, $H$ and $P$, respectively. Then the cotangent bundle $T^{*}X_{o}$ can be canonically identified with $G\times_{H}\mathfrak{h}^{\perp}$. The moment map $T^{*}X_{o}\rightarrow\mathfrak{g}^{*}$ can be written as
$\Phi_{X_{o}}^{G}\colon T^{*}X_{o}=G\times_{H}\mathfrak{h}^{\perp}\longrightarrow\mathfrak{g}^{*},\quad(g,\alpha)\longmapsto\operatorname{Ad}_{g}(\alpha)$
and it factors as
$\Phi_{X_{o}}^{G}\colon T^{*}X_{o}=G\times_{H}\mathfrak{h}^{\perp}\stackrel{{\scriptstyle q}}{{\longrightarrow}}G\times_{P}\mathfrak{h}^{\perp}\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}\mathcal{M}_{X}^{G}\subset\mathfrak{g}^{*}.$ (5.1.1)
The morphism $\psi$ is known to be generically finite and proper [Ric74] (see also [Tim11, Theorem 8.7]). We denote by $G\times_{P}\mathfrak{h}^{\perp}\rightarrow\widetilde{\mathcal{M}}_{X}\rightarrow\mathcal{M}_{X}^{G}$ the Stein factorization of $\psi$. Denote the homogeneous vector bundle $G\times_{P}\mathfrak{h}^{\perp}$ over $W$ by $E$. Then we have ~$\widetilde{\mathcal{M}}_{X}\cong\operatorname{Spec}R(\scr{O}_{\mathbb{P}E^{*}}(1))$.
5.2. Cox rings
Let $X$ be an $n$-dimensional smooth projective $G$-horospherical variety with the open orbit $X_{o}=G/H$ and $X_{o}\rightarrow W\coloneqq G/P$ the induced torus bundle. we will investigate the Cox rings of $X$ and $\mathbb{P}T_{X}$ using Theorem 4.1. For simplicity, we denote by $\mathbb{T}^{r}$ the torus $P/H$, where $r$ is the rank of $X$.
5.2.1. The Cox ring of $X$
The following result can be directly derived from [Bri07, Theorem 4.3.2] and [Gag14, Theorem 3.8]. We give an alternative proof by using Theorem 4.1.
Proposition 5.2.
Let $X$ be a smooth projective $G$-horospherical variety. Then we have
$\operatorname{Cox}(X)\cong\operatorname{Cox}(W)[S_{1},\dots,S_{m}],$
where $S_{i}$ corresponds to the canonical section $1_{D_{i}}$ for each irreducible boundary divisor $D_{i}$ of $X$. In particular, the variety $X$ is of Fano type.
Proof.
By Proposition 5.1, the torus $\mathbb{T}^{r}$ acts on $X$ with $X^{\circ}=X_{o}$ (see Theorem 4.1) and $X_{o}/\mathbb{T}^{r}=W$. Thus the description of $\operatorname{Cox}(X)$ follows from Theorem 4.1. On the other hand, since $W$ is a Fano mnaifold, the affine variety $\operatorname{Spec}\operatorname{Cox}(W)$ has klt singularities by Theorem 2.17. Then the affine variety
$\operatorname{Spec}\operatorname{Cox}(X)\cong\operatorname{Spec}\operatorname{Cox}(W)\times\mathbb{A}^{m}$
also has only klt singularities and hence $X$ is of Fano type by Theorem 2.17. ∎
Remark 5.3.
It was proved by A. Vezier in [Vez23, Corollary 2.4.2] that all the $\mathbb{Q}$-factorial projective spherical varieties are of Fano type.
5.2.2. The Cox ring of $\mathbb{P}T_{X}$
We will relate the Cox ring of $\mathbb{P}T_{X}$ to $\operatorname{Cox}(\mathbb{P}E^{*})$.
Lemma 5.4.
The dual bundle $E^{*}$ is big and globally generated. In particular, the projective bundle $\mathbb{P}E^{*}$ is a weak Fano manifold, i.e., $-K_{\mathbb{P}E^{*}}$ is big and nef.
Proof.
Note that $\mathfrak{p}^{\perp}\subset\mathfrak{h}^{\perp}$ is an invariant subspace under $P$ and $P$ acts trivially on the $r$-dimensional quotient $\mathfrak{h}^{\perp}/\mathfrak{p}^{\perp}$. Therefore we have an exact sequence
$0\longrightarrow T^{*}_{W}\cong G\times_{P}\mathfrak{p}^{\perp}\longrightarrow G\times_{P}\mathfrak{h}^{\perp}=E\longrightarrow G\times_{P}(\mathfrak{h}^{\perp}/\mathfrak{p}^{\perp})\cong\scr{O}_{W}^{\oplus r}\longrightarrow 0$
Taking dual yields
$0\longrightarrow\scr{O}_{W}^{\oplus r}\longrightarrow E^{*}\longrightarrow T_{W}\longrightarrow 0$ (5.4.1)
As $H^{1}(W,\scr{O}_{W})=0$ and $T_{W}$ is globally generated, the vector bundle $E^{*}$ is also globally generated.
On the other hand, by (5.4.1), the subvariety $\mathbb{P}T_{W}\subset\mathbb{P}E^{*}$ is a complete intersection of $r$ general elements in the sub-linear system of $|\scr{O}_{\mathbb{P}E^{*}}(1)|$ corresponding to the subspace
$H^{0}(W,\scr{O}_{W}^{\oplus r})\subset H^{0}(W,E^{*})\cong H^{0}(\mathbb{P}E^{*},\scr{O}_{\mathbb{P}E^{*}}(1)).$
So we have
$\deg(\scr{O}_{\mathbb{P}E^{*}}(1))=\deg(\scr{O}_{\mathbb{P}E^{*}}(1)|_{\mathbb{P}T_{W}})=\deg(\scr{O}_{\mathbb{P}T_{W}}(1)).$
since $T_{W}$ is big and nef, we get $\deg(\scr{O}_{\mathbb{P}E^{*}}(1))>0$. Hence $\scr{O}_{\mathbb{P}E^{*}}(1)$ is big and $\mathbb{P}E^{*}$ is a weak Fano manifold as $\scr{O}_{\mathbb{P}E^{*}}(-K_{\mathbb{P}E^{*}})\cong\scr{O}_{\mathbb{P}E^{*}}(n)$. ∎
Lemma 5.5.
Let $D$ be a prime divisor on $\mathbb{P}T_{X}$. Then the generic isotropy group of $D$ under the $\mathbb{T}^{r}$-action is trivial.
Proof.
Note that the map $f$ in Proposition 5.1 is $\mathbb{T}^{r}$-equivariant. Thus we may replace $X$ by $G\times_{P}Y$. Moreover, since the statement is clear if $D\cap\mathbb{P}T_{X_{o}}\not=\emptyset$, we may assume that $D^{\prime}\coloneqq\widebar{\pi}(D)$ is an irreducible boundary divisor contained in $X\setminus X_{o}$. Then there exists an irreducible toric divisor $F$ in $Y$ such that $D^{\prime}=G\times_{P}F$ and the isotropy subgroup $\mathbb{T}_{x}$ of the $\mathbb{T}^{r}$-action at a general point $x\in D^{\prime}$ is a $1$-dimensional torus as $F$ is of codimension one. Moreover, the induced action of $\mathbb{T}_{x}$ on $T_{x}X$ yields a decomposition $T_{D^{\prime},x}\oplus\mathbb{C}$, where $\mathbb{T}_{x}$ acts trivially on the first factor $T_{D^{\prime},x}$ and acts with weight $\in\{-1,1\}$ on the second factor. So the isotropy subgroup of a point $y\in\mathbb{P}T_{X,x}\setminus\mathbb{P}T_{D^{\prime},x}\subset D$ is trivial and we are done. ∎
Lemma 5.6.
The variety $\mathbb{P}E^{*}$ is a separation of the orbit space of the $\mathbb{T}^{r}$-action on $\mathbb{P}T_{X}$.
Proof.
Note that the map $f$ in Proposition 5.1 $\mathbb{T}^{r}$-equivariant. Thus we may replace $X$ by $G\times_{P}Y$ and consider the rational map $g\colon\mathbb{P}T_{X}\dashrightarrow\mathbb{P}E^{*}$ induced by the projectivization $\mathbb{P}T_{X_{o}}\rightarrow\mathbb{P}E^{*}$ of $q\colon T^{*}X_{o}\rightarrow E$. It is enough to show that for any prime divisor $D$ contained in $\mathbb{P}T_{X}\setminus\mathbb{P}T_{X_{o}}$, the general fiber of $D\dashrightarrow g(D)$ is exactly a $\mathbb{T}^{r}$-orbit. Let $F$ be the irreducible toric divisor in $Y$ such that $D^{\prime}=G\times_{P}F$, where $D^{\prime}=\widebar{\pi}(D)$ is a prime divisor in $X$.
Let $\mathbb{T}_{x}(\cong\mathbb{T})$ be the generic isotropy subgroup of the $\mathbb{T}^{r}$-action on $D^{\prime}$ and let $x\in D^{\prime}$ be a general point contained in the fiber of $D^{\prime}=G\times_{P}F\rightarrow W$ over the point $[P]\in W$. Then the isotropy subgroup $G_{x}$ of the $G$-action on $D^{\prime}$ at $x$ is the kernel of the following composition
$P\longrightarrow P/H\cong\mathbb{T}^{r}\longrightarrow\mathbb{T}^{r}/\mathbb{T}_{x}.$
So $P\subset N_{G}(G_{x})$ and $H\subset G_{x}$. Denote by $D^{\prime}_{o}\subset D^{\prime}$ the unique $G$-open orbit. Then the $\mathbb{T}_{x}$-action on $T^{*}X|_{D^{\prime}_{o}}$ induces a decomposition $T_{X}|_{D^{\prime}_{o}}=T_{D^{\prime}_{o}}\oplus\mathbb{C}_{D^{\prime}_{o}}$, where $\mathbb{C}_{D^{\prime}_{o}}$ is rank one trivial bundle, such that $\mathbb{T}_{x}$ acts trivially on the first factor and acts with weight $\in\{1,-1\}$ on the second factor. In particular, the general fiber of the natural projection
$(T^{*}X\setminus\textbf{0}_{X})|_{D^{\prime}_{o}}\longrightarrow T^{*}D^{\prime}_{o}$
is exactly a $\mathbb{T}_{x}$-orbit. Note that the restriction of the moment map $\varphi_{X}$ to $D^{\prime}_{o}$ factors through the following map
$T^{*}D^{\prime}_{o}\cong G\times_{G_{x}}\mathfrak{g}_{x}^{\perp}\longrightarrow G\times_{P}\mathfrak{g}_{x}^{\perp}\subset G\times_{P}\mathfrak{h}^{\perp}=E$
whose fibers are exactly the $(\mathbb{T}^{r}/\mathbb{T}_{x})$-orbits. As a consequence, the fiber of the composition
$\varphi_{X}\colon(T^{*}X\setminus\textbf{0}_{X})|_{D^{\prime}_{o}}\longrightarrow T^{*}D^{\prime}_{o}\longrightarrow G\times_{P}\mathfrak{g}_{x}^{\perp}\subset E$
is a $\mathbb{T}^{r}$-orbit and hence the general fiber of $D\dashrightarrow g(D)$ is also a $\mathbb{T}^{r}$-orbit. ∎
Now we are in the position to finish the proof of Theorem 1.10.
Proof of Theorem 1.10.
By Proposition 5.2, it only remains to consider $\mathbb{P}T_{X}$. By Theorem 2.17 and Lemma 5.4, the Cox ring $\operatorname{Cox}(\mathbb{P}E^{*})$ is finitely generated and $\operatorname{Spec}\operatorname{Cox}(\mathbb{P}E^{*})$ has only klt singularities. In particular, by Theorem 4.1 and Lemma 5.6, the Cox ring $\operatorname{Cox}(\mathbb{P}T_{X})$ is also finitely generated and hence $\mathbb{P}T_{X}$ is a Mori dream space. If additionally $X$ does not contain any $G$-stable divisor, namely $\operatorname{codim}(X\setminus X_{o})\geq 2$, then Theorem 4.1 implies that $\operatorname{Cox}(\mathbb{P}T_{X})\cong\operatorname{Cox}(\mathbb{P}E^{*})$ and hence $\mathbb{P}T_{X}$ is of Fano type by Theorem 2.17. ∎
In general it is difficult to describe the geometry of $\mathcal{Z}_{X}$ explicitly. However, since $\mathcal{Z}_{X}$ admits a natural $\mathbb{T}^{r}$-action, we have the following description of the quotient of $\mathcal{Z}_{X}$ by $\mathbb{T}^{r}$.
Proposition 5.7.
Let $X$ be a smooth projective $G$-horospherical variety without $G$-stable divisors. Then ~$\widetilde{\mathcal{M}}_{X}\cong\mathcal{Z}_{X}/\!\!/\mathbb{T}^{r}$.
Proof.
It remains to prove $S(X)^{\mathbb{T}^{r}}\cong R(\scr{O}_{\mathbb{P}E^{*}}(1))$. As $X$ does not contain $G$-stable divisors, we have $\operatorname{codim}(X\setminus X_{o})\geq 2$ and thus the natural restriction $S(X)\rightarrow S(X_{o})$ is an isomorphism of graded $\mathbb{C}$-algebras. As $T^{*}X_{o}/\mathbb{T}^{r}=E$, we obtain
$S(X_{o})^{\mathbb{T}^{r}}\cong H^{0}(T^{*}X_{o},\scr{O}_{T^{*}X_{o}})^{\mathbb{T}^{r}}\cong H^{0}(E,\scr{O}_{E})\cong R(\scr{O}_{\mathbb{P}E^{*}}(1)),$
where the second isomorphism is given by the pull back of $q\colon T^{*}X_{o}\rightarrow E$. ∎
Question 5.8.
The following two questions are inspired by Question 1.4 and Theorem 1.10.
  1. (5.8.1)
    Is $\mathbb{P}T_{X}$ of Fano type for any smooth projective horospherical variety $X$?
  2. (5.8.2)
    Find out all smooth projective horospherical varieties $X$ such that $\mathcal{Z}_{X}$ admits a symplectic resolution.
5.3. Picard number one case
In the following we focus on smooth projective $G$-horospherical varietes with Picard number one, which are completely classified by B. Pasquier. Without loss of generality, we shall assume that $X$ is not homogeneous and then there are exactly five different types ([Pas09, Theorem 0.1]):
  • $(B_{m},\omega_{m-1},\omega_{m})$ $(m\geq 3)$;
  • $(B_{3},\omega_{1},\omega_{3})$;
  • $(C_{m},\omega_{i},\omega_{i+1})$ $(m\geq 2,1\leq i\leq m-1)$;
  • $(F_{4},\omega_{2},\omega_{3})$;
  • $(G_{2},\omega_{2},\omega_{1})$.
We refer the reader to [Pas09] for the notations and we give a brief introduction to the geometry of $X$. There are exactly three $G$-orbits in $X$: the unique open $G$-orbit $X_{o}=G/H$ and two closed $G$-orbits $Z=G/P_{Z}$ and $Y=G/P_{Y}$, where $P_{Y},P_{Z}$ are two parabolic subgroups of $G$. Moreover, both $X$ and $Y$ have codimension at least two in $X$, so $\mathcal{Z}_{X}$ is symplectic by Theorem 1.6 and Theorem 1.10. The open subset $X_{Y}\coloneqq X\setminus Z$ is isomorphic to a homogeneous vector bundle
$\phi_{Y}\colon X_{Y}\cong G\times_{P_{Y}}V_{Y}\longrightarrow Y=G/P_{Y},$
where $V_{Y}$ is a simple $P_{Y}$-module, $Y\subset X_{Y}$ is just the zero section of $\phi_{Y}$ and the open subset $X_{o}$ can be identified to $X_{Y}\setminus Y$, which satisfies the following commutative diagram:
[…]
The torus $\mathbb{T}\coloneqq P/H$ has rank one and it acts on the fibers of $\phi_{Y}$ by scaling. Since $\phi_{Y}\colon X_{Y}\rightarrow Y$ can be naturally identified to the normal bundle of $Y$, the vector group $\mathbb{V}\coloneqq H^{0}(Y,N_{Y/X})$ is exactly the space of sections of $\phi_{Y}$ and it acts naturally on $X_{Y}$ as follows:
$s\cdot(g,v)\longmapsto(g,v+s(g)),$
where $s\in\mathbb{V}$. This action extends to the whole $X$ as $\operatorname{codim}Z\geq 2$. Denote by $\widetilde{G}$ the automorphism group of $X$ with $\widetilde{\mathfrak{g}}$ its Lie algebra. By [Pas09, Theorem 1.11], we have a natural isomorphism
$\widetilde{\mathfrak{g}}\cong\mathfrak{g}\oplus\mathfrak{t}\oplus\mathfrak{v},$
where $\mathfrak{t}$ and $\mathfrak{v}$ are the Lie algebras of $\mathbb{T}$ and $\mathbb{V}$, respectively. Let $\widebar{G}$ be the subgroup $G\times\mathbb{T}$ of $\widetilde{G}$. Then $X_{o}$ is also the unique open $\widebar{G}$-orbit. Let $\widebar{H}$ be the isotropic subgroup of $\widebar{G}$ at $[H]\in X_{o}$. Then $P\subset N_{\widebar{G}}(\widebar{P})$ and the natural projection to the $\mathfrak{g}$-factor induces an isomorphism $\widebar{\mathfrak{h}}^{\perp}\rightarrow\mathfrak{h}^{\perp}$ such that
$T^{*}X_{o}\cong\widebar{G}\times_{\widebar{H}}\widebar{\mathfrak{h}}^{\perp}\cong G\times_{H}\widebar{\mathfrak{h}}^{\perp}.$
So the moment map $\Phi_{X}^{\widebar{G}}\colon T^{*}X_{o}\rightarrow\mathcal{M}_{X}^{\widebar{G}}$ factors as
$T^{*}X_{o}=G\times_{H}\widebar{\mathfrak{h}}^{\perp}\longrightarrow E\cong G\times_{P}\widebar{\mathfrak{h}}^{\perp}\stackrel{{\scriptstyle h}}{{\longrightarrow}}\mathcal{M}_{X}^{\widebar{G}}\subset\widebar{\mathfrak{g}}^{*}.$
The map $h$ is generically finite and proper such that its Stein factorization coincides with $E\rightarrow\widetilde{\mathcal{M}}_{X}$. So we have the following commutative diagram:
[…]
Lemma 5.9.
The map $\Phi_{X}^{\widetilde{G}}$ is generically finite. Moreover, if $\Phi_{W}^{G}$ is birational onto its image, then the map $\Phi_{X}^{\widetilde{G}}$ is also birational.
Proof.
As $h^{-1}(\mathcal{M}_{X}^{\widebar{G}}\cap\mathfrak{g}^{*})=T^{*}W$ and the restriction $h|_{T^{*}W}\colon T^{*}W\rightarrow\mathfrak{g}^{*}$ coincides with $\Phi_{W}^{G}$, it suffices to show that a general fiber of $q$ is mapped birationally onto its image in $\mathcal{M}_{X}^{\widetilde{G}}$ under $\Phi_{X}^{\widetilde{G}}$. Fix a general fiber of $\phi_{Y}$, which is isomorphic to the vector space $V_{Y}$. Note that $\mathbb{T}$ acts on $V_{Y}$ by multiplication and the evaluation map induces a quotient $\mathbb{V}\rightarrow V_{Y}$ such that $\mathbb{V}$ acts on $V_{Y}$ via the additive action of $V_{Y}$. In particular, we have the moment map of $\mathbb{V}$ restricted to $V_{Y}$ as follows:
$\Phi_{X}^{\mathbb{V}}|_{V_{Y}}\colon T^{*}X|_{V_{Y}}\stackrel{{\scriptstyle\nu}}{{\longrightarrow}}T^{*}V_{Y}=V_{Y}\times V^{*}_{Y}\stackrel{{\scriptstyle p_{2}}}{{\longrightarrow}}V_{Y}^{*}\subset\mathfrak{v}^{*},$
where the projection $p_{2}$ coincides with the moment $\Phi_{V_{Y}}^{\mathbb{V}}$ (cf. Example 2.12). Let $O\subset T^{*}X|_{V_{Y}}$ be a general $\mathbb{T}$-orbit. Since both $\nu$ and $p_{2}$ are $\mathbb{T}$-equivariant, the image $p_{2}\circ\nu(O)$ is again a $\mathbb{T}$-orbit and the induced morphism $O\rightarrow p_{2}\circ\nu(O)$ is an isomorphism. ∎
Remark 5.10.
Now we can compute the degree of $\Phi_{W}^{G}\colon T^{*}W\rightarrow\mathcal{M}_{W}^{G}=\widebar{\mathcal{O}}_{P}$ by using Hesselink’s result ([Hes78]), which are summarized in the following table.
$X$ $m$ $i$ $\deg(\Phi_{W}^{G})$
$(B_{m},\omega_{m-1},\omega_{m})(m\geq 3)$ even - 1
odd - 2
$(B_{3},\omega_{1},\omega_{3})$ 2
$(C_{m},\omega_{i},\omega_{i+1})(m\geq 2,1\leq i\leq m-1)$ $i=m-1$ 1
$\frac{2m-2}{3}\leq i\leq m-2$ 2
$i<\frac{2m-2}{3}$ even 2
odd 4
$(F_{4},\omega_{2},\omega_{3})$ 1
$(G_{2},\omega_{2},\omega_{1})$ 1
Proof of Proposition 1.11.
We have $\mathcal{M}_{X}^{\widebar{G}}\cong\mathcal{M}_{X}^{\widetilde{G}}/\!\!/\mathbb{T}$ as $\widebar{\mathfrak{g}}=\widetilde{\mathfrak{g}}^{\mathbb{T}}$. Then the result follows from Proposition 5.7 and Lemma 5.9. ∎
6. Symplectic singularity from the quintic del Pezzo threefold
In this section we consider the quintic del Pezzo threefold $X\subset\mathbb{P}^{6}$, i.e., a smooth codimension three linear section of $\operatorname{Gr}(2,5)\subset\mathbb{P}^{9}$. We denote by $\scr{O}_{X}(1)$ the restriction $\scr{O}_{\mathbb{P}^{6}}(1)|_{X}$, which is the ample generator of $\operatorname{Pic}(X)\cong\mathbb{Z}$.
6.1. Geometry of lines on $X$
We briefly recall some facts about the lines on $X$ and refer the reader to [FN89] for more details. The Fano variety of lines on $X$ is isomorphic to $\mathbb{P}^{2}$ and the universal family $u\colon\mathcal{U}\rightarrow\mathbb{P}^{2}$ is a $\mathbb{P}^{1}$-bundle such that the evaluation morphism $\nu\colon\mathcal{U}\rightarrow X$ is a degree three finite morphism. Given a line $l$ in $X$, the restriction $T_{X}|_{l}$ is isomorphic to one of the following:
$\scr{O}_{\mathbb{P}^{1}}(2)\oplus\scr{O}_{\mathbb{P}^{1}}\oplus\scr{O}_{\mathbb{P}^{1}}\quad\textup{or}\quad\scr{O}_{\mathbb{P}^{1}}(2)\oplus\scr{O}_{\mathbb{P}^{1}}(1)\oplus\scr{O}_{\mathbb{P}^{1}}(-1).$
A line $l$ is said to be of second type if $T_{X}|_{l}$ is of the second type. The subset $B$ of $\mathbb{P}^{2}$ parametrizing lines of second type is a smooth conic curve.
The surface $R\coloneqq u^{-1}(B)$ is isomorphic to $B\times\mathbb{P}^{1}$. Then $D=\nu(R)\in|\scr{O}_{X}(2)|$ is exactly the branch locus of $\nu$ and $\nu^{-1}(D)$ decomposes into $R^{\prime}\cup R$, where $R^{\prime}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$, such that the induced morphism $R^{\prime}\rightarrow\mathbb{P}^{2}$ is a double covering. Consider the following exact sequence of vector bundles
$0\longrightarrow T_{\mathcal{U}/\mathbb{P}^{2}}\longrightarrow\nu^{*}T_{X}\longrightarrow Q\longrightarrow 0$ (6.0.1)
Then $q\colon\mathbb{P}Q\rightarrow\mathcal{U}$ is a $\mathbb{P}^{1}$-bundle and $\check{\mathcal{C}}\coloneqq f(\mathbb{P}Q)$ is the total dual VMRT of $X$, where $f\colon\mathbb{P}\nu^{*}T_{X}\rightarrow\mathbb{P}T_{X}$ [HLS22, § 2.2]. In summary, we have the following commutative diagram:
[…]
The divisor $\check{\mathcal{C}}$ is dominated by minimal sections over lines parametrized by $\mathbb{P}^{2}\setminus B$. Given a line $[l]\in\mathbb{P}^{2}\setminus B$, recall that a minimal section $\widebar{l}$ over $l$ is a section of $\mathbb{P}T_{X}|_{l}\rightarrow l$ corresponding to a trivial quotient $T_{X}|_{l}\rightarrow\scr{O}_{\mathbb{P}^{1}}$.
Lemma 6.1.
Let $C\subset X$ be an irreducible curve such that $C\not\subset D$. Then either $C$ is a line of the first type or $Q|_{\widetilde{C}}$ is ample for any component $\widetilde{C}$ of $\nu^{-1}(C)$.
Proof.
As $\widetilde{C}\not\subset R$, the map $T_{\mathcal{U}}|_{\widetilde{C}}\rightarrow\nu^{*}T_{X}|_{\widetilde{C}}$ is generically surjective and so the induced map $u^{*}T_{\mathbb{P}^{2}}|_{\widetilde{C}}\rightarrow Q|_{\widetilde{C}}$ is generically surjective. As $T_{\mathbb{P}^{2}}$ is ample, either $\widetilde{C}$ is $u$-vertical or $Q|_{\widetilde{C}}$ is ample. In the former case $C$ is a line of the first type. ∎
6.2. Weak Fano model of $\mathbb{P}T_{X}$
Our starting point is the following result obtained by A. Höring and T. Peternell in [HP24] recently.
Theorem 6.2 ([HP24, Theorem 3.1]).
Let $X$ be a smooth quintic del Pezzo threefold. Then there exists an anti-flip $\psi\colon\mathbb{P}T_{X}\dashrightarrow\mathcal{Y}$ such that $\mathcal{Y}$ is a $\mathbb{Q}$-factorial weak Fano variety with terminal singularities.
This theorem was proved in [HP24] by explicitly constructing the variety $\mathcal{Y}$. Let us briefly explain their construction (see also the proof of [HP24, Proposition 4.1]). Given a line $l$ of second type in $X$, we denote by $\widetilde{l}\subset\mathbb{P}T_{X}$ the unique negative section corresponding to the quotient $T_{X}|_{l}\rightarrow\scr{O}_{\mathbb{P}^{1}}(-1)$. By [HP24, Lemma 3.4], the locus
$S\coloneqq\bigcup_{[l]\in B}\widetilde{l}\quad\subset\quad\mathbb{P}T_{X}$
is isomorphic to $\mathbb{P}^{1}\times\mathbb{P}^{1}$. Denote by $\mu_{1}\colon\mathcal{X}_{1}\rightarrow\mathbb{P}T_{X}$ the blow-up along $S$ with exceptional divisor $E_{1}$. Then there exists a section $S_{1}$ of the projective bundle $E_{1}\rightarrow S$ ([HP24, (22) and Lemma 3.10]). Denote by $\mu_{2}\colon\mathcal{X}_{2}\rightarrow\mathcal{X}_{1}$ the blow-up along $S_{1}$ with exceptional divisor $E_{2}$. Thanks to [HP24, Lemma 3.11, Proposition 3.12 and Proposition 3.13], there exists a contraction $\nu_{2}\colon\mathcal{X}_{2}\rightarrow\mathcal{Y}_{1}$ to a projective manifold $\mathcal{Y}_{1}$, which is the blow-up along a smooth center $W_{1}$ of codimension two with $E_{2}$ being its exceptional divisor. Let $\widetilde{E}_{1}$ be the strict transform of $E_{1}$ in $\mathcal{Y}_{1}$. Then there exists a $\mathbb{P}^{2}$-bundle structure $\widetilde{E}_{1}\rightarrow S^{\prime}\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$ and [HP24, Lemma 3.15 and Proposition 4.1] imply that it can be extended to a birational contraction $\nu_{1}\colon\mathcal{Y}_{1}\rightarrow\mathcal{Y}$ such that
  • $\mathcal{Y}$ has only $\mathbb{Q}$-factorial terminal singularities, and
  • $-K_{\mathcal{Y}}$ is nef and big, and
  • the exceptional locus of $\nu_{1}$ is $\widetilde{E}_{1}$.
Denote by $W^{\prime}$ the image $\nu_{1}(W_{1})$. By [HP24, Lemma 3.3 and Proposition 3.16], we have the following commutative diagram:
[…]
such that
  • the morphism $\varphi^{+}$ is a small birational contraction with exceptional locus $S$ such that $\varphi^{+}|_{S}\colon S\cong\mathbb{P}^{1}\times\mathbb{P}^{1}\rightarrow T=\mathbb{P}^{1}$ is one of the natural projections such that its fibers are the negative sections $\widetilde{l}$ over the lines of second type and $K_{\mathbb{P}T_{X}}$ is $\varphi^{+}$-ample, and
  • the morphism $\varphi^{-}$ is a small birational contraction with exceptional locus $S^{\prime}\cup W^{\prime}$ such that $\varphi^{-1}|_{S^{\prime}}\colon S^{\prime}=\mathbb{P}^{1}\times\mathbb{P}^{1}\rightarrow T=\mathbb{P}^{1}$ is one of the natural projections, $\varphi^{-}|_{W^{\prime}}\colon W^{\prime}\rightarrow T$ is $\mathbb{P}^{2}$-bundle and $-K_{\mathcal{Y}}$ is $\varphi^{-}$-ample.
6.3. Anti-canonical model of $\mathbb{P}T_{X}$
Now we aim to describe the anti-canonical model $\mathbb{P}(\mathcal{C}_{X})$ of $\mathbb{P}T_{X}$. Let $L$ be a Weil divisor on $\mathcal{Y}$ such that $\scr{O}_{\mathbb{P}T_{X}}(\psi^{*}L)\cong\scr{O}_{\mathbb{P}T_{X}}(1)$. Then $-K_{\mathcal{Y}}=3L$. By the Basepoint Free Theorem, the $\mathbb{Q}$-Cartier divisor $L$ is semi-ample, so it defines a birational contraction
$\Phi_{mL}\colon\mathcal{Y}\longrightarrow\mathcal{Y}^{\textup{can}}.$
onto a normal variety for some $m\gg 1$. Then $-K_{\mathcal{Y}^{\textup{can}}}$ is ample and $\Phi_{mL}$ is crepant. Moreover, since $\psi$ is an isomorphism in codimension two, one can easily derive
$\mathbb{P}(\mathcal{C}_{X})\cong\operatorname{Proj}R(-K_{\mathbb{P}T_{X}})\cong\operatorname{Proj}R(-K_{\mathcal{Y}})\cong\operatorname{Proj}R(L)\cong\mathcal{Y}^{\textup{can}}.$
Denote by $\check{\mathcal{C}}^{\prime}$ the strict transform of $\check{\mathcal{C}}$ in $\mathcal{Y}$ and $H^{\prime}$ the strict transform of $\widebar{\pi}^{*}A$ for some $A\in|\scr{O}_{X}(1)|$.
Lemma 6.3.
Let $\widebar{l}$ be a general minimal section over a general line $[l]\in\mathbb{P}^{2}\setminus B$. Then $\widebar{l}\cap S=\emptyset$. In particular, the divisor $\check{\mathcal{C}}^{\prime}$ is contained in $\operatorname{Exc}(\Phi_{mL})$.
Proof.
As $D\in|\scr{O}_{X}(2)|$, the intersection $l\cap D$ consists of two points, namely $x_{1}$ and $x_{2}$. On the other hand, as $S\rightarrow D$ is one-to-one ([HP24, Lemma 3.2]), there exists a unique point $\widebar{x}_{i}\in S$ such that $\widebar{\pi}(\widebar{x}_{i})=x_{i}$ for each $i=1,2$. Note that the minimal sections over $l$ are exactly the fibers of $\mathbb{P}(N_{l/T_{X}})\cong\mathbb{P}^{1}\times l\rightarrow l$, so a general minimal section over $l$ is disjoint from the $\widebar{x}_{i}$’s and hence also $S$. Let $\widebar{l}^{\prime}$ be the strict transform of $\widebar{l}$ in $\mathcal{Y}$. As $\widebar{l}\cap S=\emptyset$, we must have
$0=-K_{\mathbb{P}T_{X}}\cdot\widebar{l}=-K_{\mathcal{Y}}\cdot\widebar{l}^{\prime}=3L\cdot\widebar{l}^{\prime}.$
Since $\check{\mathcal{C}}^{\prime}$ is dominated by such $\widebar{l}^{\prime}$, it follows that $\check{\mathcal{C}^{\prime}}$ is contained in $\operatorname{Exc}(\Phi_{mL})$. ∎
Lemma 6.4.
$\operatorname{Exc}(\Phi_{mL})=\check{\mathcal{C}}^{\prime}$.
Proof.
By [HLS22, Theorem 5.4], we have $\check{\mathcal{C}}\in|\scr{O}_{\mathbb{P}T_{X}}(3)\otimes\widebar{\pi}^{*}\scr{O}_{X}(-1)|$. On the other hand, as $\rho(\mathcal{Y})=2$, any irreducible curve $C$ contracted by $\Phi_{mL}$ is numerically proportional to $\widebar{l}^{\prime}$ by Lemma 6.3, where $\widebar{l}^{\prime}$ is the strict transform of a general minimal section $\widebar{l}$ over a general line $[l]\in\mathbb{P}^{2}\setminus B$. Then we have
$\check{\mathcal{C}}^{\prime}\cdot C=a\check{\mathcal{C}}^{\prime}\cdot\widebar{l}^{\prime}=a\check{\mathcal{C}}\cdot\widebar{l}=-aA\cdot l<0,$
where $a\in\mathbb{Q}_{>0}$. So $C$ is contained $\mathcal{C}^{\prime}$ and we are done. ∎
Lemma 6.5.
$\dim(\Phi_{mL}(\check{\mathcal{C}}^{\prime}))=3$.
Proof.
Let $F$ be an irreducible component of a general fiber of $\check{\mathcal{C}}^{\prime}\rightarrow\Phi_{mL}(\check{\mathcal{C}}^{\prime})$. We want to show that $F$ is actually a curve. Assume to the contrary that $\dim(F)\geq 2$.
According to the Step 2 and Step 3 in the proof of [HP24, Proposition 4.1], the restrictions $-K_{\mathcal{Y}}|_{S^{\prime}}$ and $-K_{\mathcal{Y}}|_{W^{\prime}}$ are ample. So $\dim(F\cap\operatorname{Exc}(\varphi^{-1}))\leq 0$. In particular, let $C$ be a general complete intersection curve contained in $F$. Then $C$ is disjoint from $\operatorname{Exc}(\varphi^{-1})$. By abuse of notation, we denote its strict transform in $\mathbb{P}T_{X}$ again by $C$. Then $C\subset\check{\mathcal{C}}$ is disjoint from $S=\operatorname{Exc}(\varphi^{+})$ and satisfies
$c_{1}(\scr{O}_{\mathbb{P}T_{X}}(1))\cdot C=L\cdot C=0.$
So $C$ is $\widebar{\pi}$-horizontal. Denote the projection $\widebar{\pi}(C)$ by $C^{\prime}$. Since $\check{\mathcal{C}}\rightarrow X$ is surjective and $F$ is general, we can assume that $C^{\prime}\not\subset R$.
Let $\widetilde{C}^{\prime}$ be an irreducible component of $\nu^{-1}(C^{\prime})$. Then there exists an irreducible curve $\widetilde{C}\subset\mathbb{P}Q$ such that $f(\widetilde{C})=C$ and $q(\widetilde{C})=\widetilde{C}^{\prime}$. As $f^{*}\scr{O}_{\mathbb{P}T_{X}}(1)|_{\mathbb{P}Q}\cong\scr{O}_{\mathbb{P}Q}(1)$, we obtain
$c_{1}(\scr{O}_{\mathbb{P}Q}(1))\cdot\widetilde{C}=0,$
therefore $Q|_{\widetilde{C}}$ is not ample. Then Lemma 6.1 implies that $C^{\prime}$ is a line of the first type and therefore $C$ is a minimal section.
Finally, since there are only three lines passing through a general point of $X$, it follows that there are also only three minimal sections passing through a general point of $\check{\mathcal{C}}$, which is a contradiction as $\dim(F)\geq 2$ and the deformations of $C$ fixing a given general point form a positive-dimensional family. ∎
For a vector bundle/vector space $E$, we denote by $\mathbf{P}E$ the geometric projectivization, i.e., $\mathbf{P}E=\mathbb{P}E^{*}$. Let $l$ be a line of the first type. Then $N_{l/X}\cong\scr{O}_{\mathbb{P}^{1}}\oplus\scr{O}_{\mathbb{P}^{1}}$ and we have the following composition of projection maps
$\Psi_{l}\colon\mathbf{P}T_{X}|_{l}\dashrightarrow\mathbf{P}N_{l/X}\cong l\times\mathbf{P}V\rightarrow\mathbf{P}V\cong\mathbb{P}^{1},$
where the first map is the projection and $V$ is a $2$-dimensional vector space.
Lemma 6.6 ([HL23, § 5.3.1]).
Let $l$ be a general line of the first type, and let $\mathcal{C}_{l}\subset\mathbf{P}T_{X}|_{l}$ be the tangent direction along $l$ of the lines meeting $l$. Then $\mathcal{C}_{l}$ is a rational curve, which is disjoint from $\mathbf{P}T_{l}$, such that $\Psi_{l}\colon\mathcal{C}_{l}\rightarrow\mathbf{P}V$ is an isomorphism.
Proposition 6.7.
The general fiber of $\check{\mathcal{C}}^{\prime}\rightarrow\Phi_{mL}(\check{\mathcal{C}}^{\prime})$ is a union of two smooth rational curves meeting at a point, which are the strict transforms of minimal sections.
Proof.
By Lemma 6.5, the general fiber $F$ of $h\colon\check{\mathcal{C}}^{\prime}\rightarrow\Phi_{mL}(\check{\mathcal{C}}^{\prime})$ is $1$-dimensional. Thus $F$ of $h$ is disjoint from $\operatorname{Exc}(\varphi^{-})$ and consists of the strict transform of minimal sections over lines of the first type. Thus it remains need to show that for a general minimal section $\widebar{l}$ over a general line $l$ in $X$, there exists exactly one other minimal section over lines meeting $\widebar{l}$. Recall that $\widebar{l}$ is a fiber of
$\mathbb{P}N_{l/X}\cong\mathbf{P}N^{*}_{l/X}=l\times\mathbf{P}V^{*}\longrightarrow\mathbf{P}V^{*},$
which corresponds a $1$-dimensional subspace $K_{\widebar{l}}$ of $V^{*}$, i.e., $\widebar{l}=l\times\{[K_{\widebar{l}}]\}$.
Let $l^{\prime}$ be another line meeting $l$ at $x\in l$ such that there exists a minimal section $\widebar{l}^{\prime}$ over $l^{\prime}$ with $\widebar{l}\cap\widebar{l^{\prime}}\not=\emptyset$. Denote by $T_{x}\subset T_{X,x}$ the $2$-dimensional subspace generated by $T_{l,x}$ and $T_{l^{\prime},x}$. Then we must have
$y\coloneqq\widebar{l}\cap\widebar{l}^{\prime}=[T_{x}]\in\mathbb{P}T_{X,x}\subset\mathbb{P}T_{X}.$
In other words, we have $y=[K_{x}]\in\mathbf{P}T^{*}_{X,x}=\mathbb{P}T_{X,x}$, where $K_{x}$ is the kernel of $T^{*}_{X,x}\rightarrow T^{*}_{x}$. Consider the induced exact sequence
$0\longrightarrow K_{x}\rightarrow N^{*}_{l/X,x}\longrightarrow T^{*}_{l^{\prime},x}\longrightarrow 0$
Under the canonical identification $N_{l/X,x}=V$, we must have $K_{x}=K_{\widebar{l}}$ and the image $\Psi_{l}([T_{l^{\prime},x}])\in\mathbf{P}V$ corresponds to the kernel $V\rightarrow K^{*}_{\widebar{l}}=K_{x}^{*}$. Then Lemma 6.6 says that such $l^{\prime}$ exists and is unique. ∎
6.4. $\mathbb{Q}$-factorial terminalization of $\mathcal{Z}_{X}$
In this subsection we want to describe the singular locus of $\mathcal{Z}_{X}$ and also its $\mathbb{Q}$-factorial terminalizations. Recall that we have the following commutative diagram:
[…]
Lemma 6.8.
The singular locus of $\mathcal{Y}$ is $S^{\prime}$ and the germ of $\mathcal{Y}$ at any point $y\in S^{\prime}$ is analytically isomorphic to
$\mathbb{C}^{3}/\mathbb{Z}_{2}\times\mathbb{C}^{2},$
where $\mathbb{Z}_{2}$ is the cyclic group generated by the action
$(x_{1},x_{2},x_{3})\mapsto(-x_{1},-x_{2},-x_{3}).$
Proof.
According to the construction in § 6.2, the singular locus of $\mathcal{Y}$ is contained in $S^{\prime}=\nu_{1}(\widetilde{E}_{1})$ as $\mathcal{Y}_{1}$ is smooth and $\widetilde{E}_{1}\rightarrow S^{\prime}$ is a $\mathbb{P}^{2}$ bundle such that the restriction of $\scr{O}_{\mathcal{Y}_{1}}(\widetilde{E}_{1})$ to any fiber is isomorphic to $\scr{O}_{\mathbb{P}^{2}}(-2)$ by [HP24, Proposition 3.15]. The result is obtained by cutting $\mathcal{Y}$ with $2$ general hypersurfaces passing through $y$. ∎
Lemma 6.9.
The singular locus of $\mathcal{L}$ is ${\bf 0}_{S^{\prime}}$ and the germ of $\mathcal{L}$ at any point $z\in{\bf 0}_{S^{\prime}}$ is analytically isomorphic to
$\mathbb{C}^{4}/\mathbb{Z}_{2}\times\mathbb{C}^{2},$
where ${\bf 0}_{S^{\prime}}={\bf 0}_{\mathcal{L}}\cap p^{-1}(S^{\prime})$ and $\mathbb{Z}_{2}$ is the cyclic group generated by the action
$(x_{1},x_{2},x_{3},x_{4})\mapsto(-x_{1},-x_{2},-x_{3},-x_{4}).$
In particular, the variety $\mathcal{L}$ has only $\mathbb{Q}$-factorial terminal singularities.
Proof.
As $\mathcal{Y}$ is smooth outside $S^{\prime}$, the divisor $L$ is Cartier outside $S^{\prime}$. So the singular locus of $\mathcal{L}$ is contained in $p^{-1}(S^{\prime})$. We note also that any point $y\in S^{\prime}$ has index two and $-K_{\mathcal{Y}}$ is the generator of the local divisor class group around $y$. In particular, as $-K_{\mathcal{Y}}=3L$, we have $L\sim-K_{\mathcal{Y}}$ around $y$. Then an easy computation shows that the germ of $\mathcal{L}$ over $y$ is analytically isomorphic to
$\mathbb{C}^{4}/\mathbb{Z}_{2}\times\mathbb{C}^{2},$
and its singular locus is just $\{(0,0,0,0)\}\times\mathbb{C}^{2}$; that is, it is contained in the zero section ${\bf 0}_{\mathcal{L}}=\{x_{4}=0\}$. ∎
Proposition 6.10.
The symplectic variety $\mathcal{Z}_{X}$ admits a unique $\mathbb{Q}$-factorial terminalization $\mu\colon\mathcal{Z}^{\prime}_{X}\rightarrow\mathcal{Z}_{X}$. Set $Y\coloneqq\mu^{-1}(0)$. Then the following statements hold:
  1. (6.10.1)
    The birational map $\mathcal{L}\rightarrow\mathcal{Z}_{X}$ induces an isomorphism $\mathcal{L}\setminus{\bf 0}_{\mathcal{L}}\cong\mathcal{Z}^{\prime}_{X}\setminus Y$.
  2. (6.10.2)
    The fiber $Y$ is irreducible and $\mu$ is a symplectic resolution outside 00.
  3. (6.10.3)
    $\mathcal{Z}_{X,\operatorname{sing}}=Z$ and $\mathcal{Z}_{X}$ has $A_{2}$-singularities along the general points of $Z$.
Proof.
Set $E=\textbf{0}_{\mathcal{L}}$. Then $E\rightarrow X$ is an isomorphism and $\scr{O}_{E}(E)\cong\scr{O}_{X}(-L)$. On the other hand, as $\varphi_{L}^{*}K_{\mathcal{Z}_{X}}=K_{\mathcal{L}}-2E$, running a relative $K_{\mathcal{L}}$-MMP over $\mathcal{Z}_{X}$ yields a birational contraction $g\colon\mathcal{L}\dashrightarrow\mathcal{Z}^{\prime}_{X}$ to a normal variety with $\mathbb{Q}$-factorial terminal singularities such that it satisfies the following commutative diagram
[…]
where $\check{\mathcal{C}}^{\prime}_{a}=g(p^{-1}(\check{\mathcal{C}}^{\prime}))$. Note that $-E$ is $\varphi_{L}$-big and $\varphi_{L}$-nef. In particular, as $K_{\mathcal{Z}^{\prime}_{X}}=g_{*}{K_{\mathcal{L}}}$ is $\mu$-nef, the Negativity Lemma implies that $g_{*}E=0$ and thus $\operatorname{Exc}(g)=E$ as $\operatorname{Exc}(g)\subset E$. Moreover, since $\mathcal{L}$ is smooth outside $E$ and $Y=g(E)$, it follows that $Y$ is irreducible and $\mathcal{Z}^{\prime}_{X}$ is smooth outside $Y$. This proves (6.10.1) and (6.10.2).
For (6.10.3), since $\rho(\mathcal{Z}^{\prime}_{X})=1$ and $\check{\mathcal{C}}_{a}^{\prime}$ is contracted by $\mu$, the relative movable cone of $\mu$ is $\mathbb{R}_{\geq 0}[\hat{\mathcal{C}}^{\prime}]$. Hence $\mathcal{Z}^{\prime}_{X}$ is the unique $\mathbb{Q}$-factorial terminalization of $\mathcal{Z}_{X}$. Note that $\mu$ is a crepant map, which contracts $\check{\mathcal{C}}^{\prime}_{a}$ to $Z$ (cf. Lemma 6.4). In particular, since $\mathcal{Z}^{\prime}_{X}$ is smooth outside $Y$, the variety $\mathcal{Z}_{X}$ is exactly singular along $Z$ $(\ni 0)$. Then the statement follows from Proposition 6.7. ∎
Proposition 6.11.
The variety $\mathcal{Z}^{\prime}_{X}$ is not smooth. In particular, the symplectic orbifold cone $\mathcal{Z}_{X}$ does not admit any symplectic resolution.
Proof.
Assume that $\mathcal{Z}_{X}$ admits a symplectic resolution. Then $F$ is smooth and $\mathcal{Z}^{\prime}_{X}\cong T^{*}F$ by Corollary 2.11 and hence $F$ is a rational homogeneous space as $\dim(\mathcal{Z}_{X})=6$. In particular, we have
$3=\dim H^{0}(X,T_{X})=\dim S(X)_{1}=\dim S(Y)_{1}=\dim H^{0}(Y,T_{Y})>3,$
which is a contradiction. ∎
Acknowledgments
The authors thank Michel Brion and Boris Pasquier for helpful discussions on horospherical varieties. J. Liu is supported by the National Key Research and Development Program of China (No. 2021YFA1002300) and the Youth Innovation Promotion Association CAS. Both authors are supported by the CAS Project for Young Scientists in Basic Research (No. YSBR-033) and the NSFC grant (No. 12288201).